### v1

# Generalized Hausdorff dimension of irrationals with Lagrange value exactly 3

Abstract.
We study the generalized Hausdorff dimension of some natural subsets of $k^{-1}(3)$, where $k^{-1}(3)$ consists of the real numbers $x$ for which $\left|x-\frac{p}{q}\right|<\frac{1}{(3+\varepsilon)q^{2}}$ has infinitely many rational solutions $\frac{p}{q}$ for any $\varepsilon<0$ but only finitely many for any $\varepsilon>0$. It is well known that $k^{-1}(3)$ is an uncountable set with Hausdorff dimension zero. Given any dimension function $h$, we determine the exact “cut point” at which the generalized Hausdorff dimension $\mathcal{H}^{h}(k^{-1}(3))$ drops from infinity to zero. In particular we show that such a measure is always zero or not $\sigma$–finite, and, as an application, we can classify topologically $k^{-1}(3)$. Moreover, we show that the subset of attainable elements of $k^{-1}(3)$ has the same generalized Hausdorff dimension as $k^{-1}(3)$, but the subset of non–attainable elements of $k^{-1}(3)$ has a “strictly smaller” generalized Hausdorff dimension.
Key words and phrases:
Lagrange spectrum, Generalized Hausdorff measure
2020 Mathematics Subject Classification:
11J06, 11J70, 28A78.
1. Introduction
The Hausdorff dimension provides a powerful tool for quantifying the fractal structure of sets in metric spaces, but it is often too coarse to distinguish between certain extremely thin sets, all of which may have Hausdorff dimension zero. In such cases, a finer scale of measurement is required. To achieve this, one can introduce the generalized Hausdorff measure
$\mathcal{H}^{h}(E)=\lim_{\delta\to 0}\inf\left\{\sum_{i}h(\operatorname{diam}(U_{i})):E\subset\bigcup_{i}U_{i},\,\operatorname{diam}(U_{i})<\delta\right\},$
where $h:(0,\infty)\to(0,\infty)$ is an nondecreasing and right continuous function such that $h(r)\to 0$ as $r\to 0$. Such a function is called a dimension function or gauge function. See [11] for a comprehensive treatment of these (outer) measures.
For the special choice $h(r)=r^{s}$, one recovers the classical $s$-dimensional Hausdorff measure. By allowing more general gauge functions, one can distinguish between sets of Hausdorff dimension zero that still exhibit substantial size or complexity when measured at a finer scale.
This approach has proved particularly useful for exceptional sets arising in Diophantine approximation and dynamical systems (for example [1]), by allowing one to determine for which gauge functions $h$ the Hausdorff hh–measure $\mathcal{H}^{h}$ of a set of zero Hausdorff dimension is positive, infinite, or zero. A classical example is the set of Liouville numbers $\mathbb{L}$. Recall that a real number $x\in\mathbb{R}\setminus\mathbb{Q}$ is called a Liouville number if, for every $n\in\mathbb{N}$ there exist integers $p$ and $q$ with $q>1$ such that
$\left|x-\frac{p}{q}\right|<\frac{1}{q^{n}}.$
Olsen and Renfro determined the exact cut point between gauge functions $h$ for which $\mathcal{H}^{h}(\mathbb{L})$ is infinite and those for which it is zero (see [8] and [9]). Furthermore, they proved that if $\mathcal{H}^{h}(\mathbb{L})=\infty$ for a given gauge function $h$, then $\mathbb{L}$ does not have $\sigma$-finite $\mathcal{H}^{h}$ measure.
At the other extreme, for badly approximable numbers, another important class of sets arising in Diophantine approximation consists of numbers sharing the same Diophantine approximation constant. That is, for $x\in\mathbb{R}\setminus\mathbb{Q}$, the best constant of Diophantine approximation of $x$ is
$k(x)=\sup\left\{c>0\mid\left|x-\frac{p}{q}\right|<\frac{1}{cq^{2}}\text{ has infinitely many solutions }\frac{p}{q}\in\mathbb{Q}\right\}$
where $k$ can be regarded as a map from $\mathbb{R}\backslash\mathbb{Q}$ to $(0,\infty]$. The image $\{k(x):x\in\mathbb{R}\setminus\mathbb{Q},k(x)<\infty\}$ of this function is known as the Lagrange spectrum.
It is well known that $k^{-1}(3)$ is an uncountable dense set and it is a consequence of [7, Theorem 1] that $k^{-1}(3)$ has Hausdorff dimension 0. Indeed, for $t\in\mathbb{R}$ define
(1) $K_{t}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}^{*})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq t,\forall k\in\mathbb{Z}\}.$
Then $K_{t}\subset k^{-1}(-\infty,t]$ and $D(t)=\dim_{H}(K_{t})=\overline{\dim_{B}}(K_{t})=\dim_{H}(k^{-1}(-\infty,t])$ is a continuous function with $\max\{t\in\mathbb{R}:D(t)=0\}=3$.
In this paper, we determine the exact cut point at which the Hausdorff hh–measure of $k^{-1}(3)$ drops from infinity to zero. This provides a complete characterization of all Hausdorff measures $\mathcal{H}^{h}(k^{-1}(3))$ without imposing any specific regularity assumptions on the dimension function $h$. Moreover, we show that if $\mathcal{H}^{h}(k^{-1}(3))=\infty$, then $k^{-1}(3)$ does not possess $\sigma$-finite $\mathcal{H}^{h}$ measure. Our main theorem is the following:
Theorem 1.1.
Let $h$ be a dimension function.
  1. (1)
    If
    $\limsup_{\varepsilon\to\infty}\frac{\log h(\varepsilon)}{\log\varepsilon}=0$
    then the set $k^{-1}(3)$ does not have $\sigma$-finite $\mathcal{H}^{h}$ measure.
  2. (2)
    If
    $\limsup_{\varepsilon\to\infty}\frac{\log h(\varepsilon)}{\log\varepsilon}>0$
    then $\mathcal{H}^{h}(k^{-1}(3))=0$.
Moreover, if $A_{3}$ consists of those $x\in k^{-1}(3)$ which are attainable, then $\mathcal{H}^{h}(A_{3})=\mathcal{H}^{h}(k^{-1}(3))$ for any dimension function $h$.
Definition 1.2.
An irrational number $x\in k^{-1}(3)$ is called attainable if $\left|x-\frac{p}{q}\right|\leq\frac{1}{3q^{2}}$ has infinitely many rational solutions $\frac{p}{q}$.
In comparison, one can say that the set of Liouville numbers $\mathbb{L}$ is strictly larger than $k^{-1}(3)$, in the sense that any dimension function $h$ for which $\mathcal{H}^{h}(k^{-1}(3))=\infty$ also satisfies $\mathcal{H}^{h}(\mathbb{L})=\infty$, whereas the converse does not necessarily hold. We will provide an explicit example of a dimension function $h$ such that $\mathcal{H}^{h}(\mathbb{L})=\infty$ while $\mathcal{H}^{h}(k^{-1}(3))=0$.
A more general property of Liouville numbers is that they are immeasurable, that is, any translation invariant Borel measure on $\mathbb{R}$ either assigns zero measure to $\mathbb{L}$ or is not $\sigma$–finite on $\mathbb{L}$ [3, Theorem 1.1] (note that this includes the Hausdorff hh–measures). The main result of [3] applies to a broad class of sets, however, it does not apply to $k^{-1}(3)$ because of several reasons. Recall that set of Liouville numbers $\mathbb{L}$ forms a $G_{\delta}$ set that is invariant by rational translations (it is not difficult to see that $k^{-1}(3)$ is not invariant by such translations). A big difference is that from the topological point of view, the Liouville numbers are also larger than $k^{-1}(3)$.
Theorem 1.3.
The set $k^{-1}(3)$ is not $G_{\delta}$ and not $F_{\sigma}$, but it is $F_{\sigma\delta}$.
On the other hand, there are certain subsets of $k^{-1}(3)$ that are natural to consider. The use of generalized Hausdorff dimension is then necessary to distinguish the size of these subsets. For example, the set $K_{3}$ is a Cantor set with Hausdorff dimension 0, since $K_{3}\subset k^{-1}(-\infty,3]$. For $t>3$, the Cantor set $K_{t}$ has the same positive Hausdorff dimension as $k^{-1}(-\infty,t]$, so it is natural to ask whether $K_{3}$ is as large as $k^{-1}(-\infty,3]$ from the point of view of Hausdorff dimension. Since $k^{-1}(\infty,3)$ is countable because of Markov’s theorem [2, Theorem 16], it suffices to compare $K_{3}$ with $k^{-1}(3)$. The following result shows that $K_{3}$ is, in fact, significantly smaller than $k^{-1}(3)$.
Theorem 1.4.
Let $h$ be a dimension function.
  1. (1)
    If
    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>3$
    then $\mathcal{H}^{h}(K_{3})=0$.
A forthcoming work of the first two authors with Zhe Cao will imply the following result: for any $x=[x_{0};x_{1},x_{2},\dots]\in k^{-1}(3)$ such that the inequality $\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}$ has only finitely many solutions, there is an $N\in\mathbb{N}$ such that $0;x_{N},x_{N+1},\dots]\in K_{3}$. In particular, if we denote by $B_{3}$ the set of non–attainable elements of $k^{-1}(3)$, then for example for the dimension function $h(x)=\frac{1}{|\log\varepsilon|^{4}}$ one has that $\mathcal{H}^{h}(k^{-1}(3))=\infty$ but $\mathcal{H}^{h}(B_{3})=0$. In conclusion, there is a large gap between the generalized Hausdorff dimension of these sets, so the set of non–attainable elements $x\in k^{-1}(3)$ is also significantly smaller than $k^{-1}(3)$.
2. Preliminaries
Given an nondecreasing and right continuous function $h:[0,\infty)\to[0,\infty)$ with $h(0)=0$, we define the corresponding Hausdorff measure $\mathcal{H}^{h}$ as
$\mathcal{H}^{h}(E):=\lim_{\delta\to 0}\inf\left\{\sum h(\operatorname{diam}E_{i}):E_{i}\text{ open},E\subset\bigcup E_{i},\operatorname{diam}E_{i}\leq\delta\right\}.$
We say that a set $E$ is an hh–set if $0<\mathcal{H}^{h}(E)<\infty$. It is not always possible to find a dimension function for which this holds.
Given $\underline{a}\in(\mathbb{N}^{*})^{\mathbb{Z}}$ define
$\lambda_{i}((a_{n})_{n\in\mathbb{Z}})=[a_{i};a_{i+1},a_{i+2},\dots]+[0;a_{i-1},a_{i-2},\dots].$
By abuse of notation, if $(a_{n})_{n\geq 0}$ is a one-sided sequence of positive integers, we also denote
$\lambda_{i}((a_{n})_{n\geq 0})=[a_{i};a_{i+1},a_{i+2},\dots]+[0;a_{i-1},a_{i-2},\dots,a_{1}].$
Continued fractions are useful to compute $k(x)$. Indeed, it is a classical fact that if $x=[a_{0};a_{1},a_{2},\dots]$ then
(2) $\left|x-\frac{p_{n}}{q_{n}}\right|=\frac{1}{(\gamma_{n+1}+\eta_{n+1})q_{n}^{2}},$
where $\gamma_{n+1}=[a_{n+1};a_{n+2},\dots]$ and $\eta_{n+1}=[0;a_{n},a_{n-1},\dots,a_{1}]$. In particular we have the formula $k(x)=\limsup_{i\to\infty}\lambda_{i}(x)$.
A curious and crucial identity to study irrational numbers $x$ with values $k(x)$ near 3 has been the following:
(3) $2;2,z]+[0;1,1,z]=3,\qquad\text{for all $z>0$}.$
Given a finite word of positive integers $w=(a_{1},\dots,a_{n})\in(\mathbb{N}^{*})^{n}$, define the closed sub-interval of $[0,1]$
$I(w):=\{x\in[0,1]\ \mid\ x=[0;a_{1},a_{2},\dots,a_{n},x],x\geq 1\}\cup\{[0,a_{1},a_{2},\dotsc,a_{n}]\},$
consisting of the numbers in $[0,1]$ whose continued fractions start with $w$.
The precise diameter of such intervals is given by
$\operatorname{diam}I(w)=\frac{1}{q_{n}(q_{n}+q_{n-1})},$
where $q_{k}$ is defined recursively by $q_{0}=1$, $q_{1}=a_{1}$ and $q_{k+2}=a_{k+2}q_{k+1}+q_{k}$.
In particular, if $a_{1},\dots,a_{n}\in\{1,2\}$, is easy to see that
(4) $(3+2\sqrt{2})^{-n-1}<\operatorname{diam}I(w)<\left(\frac{3+\sqrt{5}}{2}\right)^{-n+1}$
Given $t>0$, define the subshift
$\Sigma_{t}=\left\{\underline{a}\in(\mathbb{N}^{*})^{\mathbb{Z}}\ \mid\ \sup_{n\in\mathbb{Z}}\lambda_{n}(\underline{a})\leq t\right\}.$
Note that for $t\in(0,\infty)$, one only needs finitely many symbols because $\Sigma_{t}\subset\{1,\dots,\lfloor t\rfloor\}^{\mathbb{Z}}$. Moreover, for $t=\sqrt{12}$ it is known that $\Sigma_{t}=\{1,2\}^{\mathbb{Z}}$.
In general, the set $K_{t}$ is always closed, but it is not necessarily a Cantor set. For instance, when $t=\sqrt{13}$ the set $K_{t}$ includes the isolated point $0;\overline{3}]$. Another examples occur for all $\sqrt{5}\leq t<3$, where the set $K_{t}$ is countable. More generally, $K_{t}$ is a Cantor set whenever $\Sigma_{t}$ is a transitive subshift. Situations where the subshift $\Sigma_{t}$ is not transitive are for example when $t$ lies at the right endpoint of a gap of the Markov spectrum. Another particular situation when $\Sigma_{t}$ is not transitive is precisely at $t=3$: the periodic orbits $\dots 1111\dots$ and $\dots 2222\dots$ are both contained in $\Sigma_{3}$, yet there is no orbit in $\Sigma_{3}$ that contains both finite subwords $1111$ and $2222$. However, the set $K_{3}$ is a Cantor set, as we will see.
The subshift $\Sigma_{3}$ was characterized in [10]: after substituting $22\mapsto a$ and $11\mapsto b$ it coincides with the subshift of all Sturmian bi-infinite words (eventually periodic and aperiodic). It is well known that the closure of any aperiodic Sturmian bi-infinite word gives a minimal infinite subshift, so in particular, any orbit in that closure is an accumulation point. On the other hand, the eventually periodic Sturmian bi-infinite words can be approximated by the aperiodic (in the shift topology), thus any orbit of $\Sigma_{3}$ is an accumulation point. In particular the shift $\Sigma_{3}$ is homeomorphic to a Cantor space and consequently the set $K_{3}$ is a Cantor subset of $\mathbb{R}$.
We define $\Sigma(t,n)$ to be the set of length-$n$ subwords of sequences in $\Sigma_{t}$. From [4, Theorem 1.1] we have the following:
Theorem 2.1.
For all $n\geq 68$ we have
$\Sigma(3+6^{-3n},n)=\Sigma(3,n)=\Sigma(3-6^{-3n},n).$
It is well known [6] that the cardinality of all Sturmian factors is asymptotically equivalent to $n^{3}/\pi^{2}$. In particular, we have a good control on the cardinality of the subfactors of $\Sigma_{3}$.
Lemma 2.2.
$|\Sigma(3,n)|\sim n^{3}/(4\pi^{2})$.
Proof.
Let $L_{n}$ denote the set of all Sturmian factors of length $n$. From [6] we know that $|L_{n}|\sim n^{3}/\pi^{2}$. We will prove the following:
$|\Sigma(3,n)|=\begin{cases}|L_{n/2}|+|L_{n/2+1}|,&\text{if $n$ is even},\\ 2|L_{(n+1)/2}|,&\text{if $n$ is odd}.\end{cases}$
First suppose $n$ even. Given $w\in\Sigma(3,n)$, if $w$ begins with an even block of 1’s or 2’s then by applying the inverse substitution $22\mapsto a$, $11\mapsto b$ it can be written as an Sturmian factor of length exactly $n/2$. If it begins with an odd block, then it also ends with an odd block so by adding one digit $\{1,2\}$ (which are uniquely determined because $w$ contains both 1 and 2 in this case) at each side and applying the substitution, we obtain a Sturmian word of length $n/2+1$.
Now assume $n$ odd. In this case the elements of $\Sigma(3,n)$ can be obtained from $L_{(n+1)/2}$ by substituting $a\mapsto 22$, $b\mapsto 11$ and then deleting precisely the first or the last digit. ∎
Remark 2.3.
In fact, it is possible to prove that $|\Sigma(3,n)|\leq 9n^{3}$ for all $n\geq 1$, see [4, Corollary 3.13].
Lemma 2.4.
Let $t\in L$ and $\varepsilon>0$. Given any $x=[0;a_{0};a_{1},a_{2},\dots]\in k^{-1}(t)\cap(0,1)$, there is a $n_{0}$ and a sequence of positive integers $(b_{j})_{j<n_{0}}$ such that
$(\dots,b_{n_{0}-2},b_{n_{0}-1},a_{n_{0}},a_{n_{0}+1},a_{n_{0}+2},\dots)\in\Sigma_{t+\varepsilon}.$
Proof.
First, note that for $n$ large $a_{n},a_{n+1},\dots\in\{1,2,\dots,\lfloor t\rfloor\}$. Since $k(x)=\limsup_{i\to\infty}\lambda_{i}(x)=t$, there is a subsequence $(i_{k})_{k}$ such that $\lambda_{k\to\infty}\lambda_{i_{k}}(x)=t$. In particular, by taking a further subsequence, we can assume that the subwords $(a_{i_{k}}),(a_{i_{k+1}-1},a_{i_{k+1}},a_{i_{k+1}+1}),(a_{i_{k+2}-2},a_{i_{k+2}-1},a_{i_{k+2}},a_{i_{k+2}+1},a_{i_{k+2}+2}),\dots$ converge to a bi-infinite sequence $\underline{b}=(b_{n})_{n\in\mathbb{Z}}$ with $\sup_{n\in\mathbb{Z}}\lambda_{i}(\underline{b})=t$. Take $m=\lceil\log_{2}\varepsilon\rceil+1$ and $n_{0}:=i_{k}$ so large such that $\lambda_{i}(x)<t+\varepsilon/2$ for all $i\geq i_{k}$ and such that $(a_{i_{k}-m},\dots,a_{i_{k}},\dots,a_{i_{k}+m})=(b_{i_{k}-m},\dots,b_{i_{k}},\dots,b_{i_{k}+m})$. In particular we have that $(\dots,b_{n_{0}-2},b_{n_{0}-1},a_{n_{0}},a_{n_{0}+1},a_{n_{0}+2},\dots)$ has Markov value at most $t+\varepsilon/2+2^{-m}<t+\varepsilon$. ∎
3. Dimensionality of $k^{-1}(3)$
Recall that $k^{-1}(3)$ denotes the set of real numbers $x\in\mathbb{R}$ such that,
$\left|x-\frac{p}{q}\right|\leq\frac{1}{(3+\varepsilon)q^{2}}$
has infinitely many rational solutions $\frac{p}{q}$ for any $\varepsilon<0$ but only finitely many for any $\varepsilon>0$. An interesting subset of these numbers is the set of attainable numbers, defined as follows.
Definition 3.1.
An irrational number $x\in k^{-1}(3)$ is called attainable if $\left|x-\frac{p}{q}\right|\leq\frac{1}{3q^{2}}$ has infinitely many rational solutions $\frac{p}{q}$.
Now we will show that there is a large class of numbers in $k^{-1}(3)$ which are attainable.
Lemma 3.2.
Let $(e_{i})_{i\geq 1}$ be any sequence of positive integers with $\lim_{i\to\infty}e_{i}=\infty$. Then the irrational number
$x=[0;1^{e_{1}},2,2,1^{e_{2}},2,2,1^{e_{3}},2,2,\dots],$
satisfies $k(x)=3$ and is attainable.
Proof.
Let $i_{0}\geq 2$ be such that $e_{i}\geq 2$ for all $i\geq i_{0}$. If $e_{i+1}$ is even and $e_{i}$ odd, or if $e_{i}<e_{i+1}+2$ are both odd, then using (2) one has
(5) $\begin{split}\lambda(\dots 221^{e_{i}}|221^{e_{i+1}}22\dots)&=[2;2,1^{e_{i+1}},2,2,\dots]+[0;1^{e_{i}},2,2,\dots]\\ &=3+[0;1^{e_{i}},2,2,\dots]-[0;1^{e_{i+1}+2},2,2,\dots]>3.\end{split}$
If $e_{i+1}$ is odd and $e_{i}$ even, or if $e_{i+1}>e_{i}+2$ are both even, or if $e_{i+1}<e_{i}+2$ are both odd, then using (2) again
(6) $\begin{split}\lambda(\dots 221^{e_{i}}2|21^{e_{i+1}}22\dots)&=[0;1^{e_{i+1}},2,2,\dots]+[2;2,1^{e_{i}},2,2,\dots]\\ &=3+[0;1^{e_{i+1}},2,2,\dots]-[0;1^{e_{i}+2},2,2,\dots]>3.\end{split}$
Note that, disregarding parities, since $e_{i}\to\infty$ as $i\to\infty$, the equations (5) and (6) show that $k(x)=\lim_{n\to\infty}\lambda_{n}(x)=3$.
Now we will show that for infinitely many positions $n$ one has $\lambda_{n}(x)>3$. First, assume that there infinitely many indices $i\geq i_{0}$ such that $e_{i}$ and $e_{i+1}$ have different parities. From both equations above we see that we have infinitely many positions $n$ such that $\lambda_{n}(x)=\gamma_{n+1}+\eta_{n+1}>3$.
Now, assume that after some larger $i_{0}$, all $e_{i}$ have the same parity for $i\geq i_{0}$. If such parity is odd, then using (5) and (6) it suffices to consider those infinitely many indices such that $e_{i}<e_{i+1}+2$ or $e_{i+1}<e_{i}+2$. If the parity is eventually even and $e_{i+1}>e_{i}+2$ infinitely many times, then using (6) we are done. Otherwise, the pattern $221^{2k}221^{2k+2}221^{2k+4}221^{2k+6}$ will repeat infinitely many times (with increasing $k$) and in this case we use (2) again to conclude that
$\lambda(\dots 221^{2k}221^{2k+2}2|21^{2k+4}221^{2k+6}\dots)$
$=[0;1^{2k+4},2,2,1^{2k+6},\dots]+[2;2,1^{2k+2},2,2,1^{2k},2,2,\dots]$
$=3+[0;1^{2k+4},2,2,1^{2k+6},\dots]-[0;1^{2k+4},2,2,1^{2k},2,2,\dots]>3,$
infinitely many times.
Dimension functions $h$ for which $\mathcal{H}^{h}(h^{-1}(3))=\infty$ will be obtained by constructing suitable Cantor subsets $C\subset k^{-1}(3)$. In view of this, we introduce the following lemma concerning the fractal geometry of Cantor sets. The next lemma is a slightly modified version of [5, Lemma 1]. The proof is similar, and we include it here for completeness.
Lemma 3.3.
Let $K\subset\mathbb{R}$ be a Cantor set obtained as $K=\bigcap_{n\geq 1}K_{n}$ where $K_{n}$ is a union of intervals of size at least $\varepsilon_{n}>0$ with disjoint interiors such that, for each interval $I$ of $K_{n}$, we have
$|\left\{J\text{ interval of }K_{n+1}|\ J\subset I\right\}|\geq m(n)>0.$
Let $N_{k}=\Pi_{n<k}m(n)$, and let $h:[0,\infty)\to[0,\infty)$ be a dimension function such that $\lim_{k\to\infty}N_{k}\cdot h(\varepsilon_{k+1})=\infty$, then $\mathcal{H}^{h}(K)=\infty$.
Proof.
Let us consider a finite covering of $K$ by intervals $K\subset\bigcup_{j=1}^{\ell}I_{j}$, and denote, for any $j\in\{1,\ldots,\ell\}$, by $v(j)$ the unique integer such that $\varepsilon_{v(j)+1}\leq|I_{j}|<\varepsilon_{v(j)}$.
We now replace each interval $I_{j}$ by at most two intervals of $K_{v(j)}$ whose union contains $I_{j}\cap K$. Note that this is possible because if $I_{j}$ intersects three or more intervals of $K_{v(j)}$, then $I_{j}$ must contain one of them, and hence $\varepsilon_{v(j)}<\operatorname{diam}(I_{j})$. So, we get a covering of $K$ by intervals of size at least $\varepsilon_{v(j)}$ of $K_{v(j)}$, from which we can extract a sub-covering
$K\subset\bigcup_{s=1}^{\tilde{\ell}}\tilde{I}_{s}$
by disjoint intervals $\tilde{I}_{s}$ of sizes at least $\varepsilon_{n(s)}$ of $K_{n(s)}$, and since each $I_{j}$ was subdivided into at most two intervals, we have, for every positive integer $n$,
(7) $|\{1\leq s\leq\tilde{\ell}\mid n(s)=n\}|\leq 2|\{1\leq j\leq\ell\mid v(j)=n\}|.$
Claim.
If we have a finite collection of disjoint intervals $\tilde{I}_{s}$ covering $K$ such that, for all $s$, $\tilde{I}_{s}$ is an interval of $K_{n(s)}$ for some $n(s)$, and if we denote $k_{n}:=|\{1\leq s\leq\tilde{\ell}\mid n(s)=n\}|$, then we have
$\sum_{n\geq 1}k_{n}/N_{n}\geq 1.$
To prove the claim, let us first remark that it is trivial when $n(s)$ is constant and equal to $\tilde{n}$, since in this case the covering $\bigcup_{s=1}^{\tilde{\ell}}\tilde{I}_{s}$ coincides with $K_{\tilde{n}}$ which in turn is the union of at least $N_{\tilde{n}}$ disjoint intervals of size $\varepsilon_{\tilde{n}}$, so $\sum_{n\geq 1}k_{n}/N_{n}=k_{\tilde{n}}/N_{\tilde{n}}\geq 1$.
Otherwise, let us take $n^{*}$ maximal, such that $k_{n^{*}}\neq 0$, and change all intervals of $K_{n^{*}}$ which belong to the collection and intersect a given interval $I$ of $K_{n^{*}-1}$ by $I$. This is possible because, whenever one interval of $K_{n^{*}}$ in the collection intersects $I$, all intervals of $K_{n^{*}}$ intersecting $I$ necessarily belong to the collection. By doing so, we replace at least $N_{n^{*}}/N_{n^{*}-1}$ intervals of $K_{n^{*}}$ by one interval of $K_{n^{*}-1}$, so that the positive integer $\sum_{n\geq 1}n\cdot k_{n}$ diminishes and, since
$(k_{n^{*}}-N_{n^{*}}/N_{n^{*}-1})/N_{n^{*}}+(k_{n^{*}-1}+1)/N_{n^{*}-1}=k_{n^{*}}/N_{n^{*}}+k_{n^{*}-1}/N_{n^{*}-1},$
the value of $\sum_{n\geq 1}k_{n}/N_{n}$ does not increase. We repeat this process until all intervals belong to the same $K_{r}$, when we have at least $N_{r}$ intervals. Since $N_{r}/N_{r}=1$, the claim is proved.
Now we can conclude the proof of Lemma 3.3.
As we have seen above, (some of) the intervals of size at least $\varepsilon_{v(j)}$ we considered form a covering of $K$ by disjoint intervals $\tilde{I}_{s}$ where, for all positive integers $1\leq s\leq\tilde{\ell}$, $\tilde{I}_{s}$ is an interval of $K_{n(s)}$ for some $n(s)$. It follows from the claim and from (7) that
$2\sum_{r\geq 1}(m_{r}/N_{r})\geq 1,$
where we put $m_{r}=|\{j\mid m(j)=r\}|$.
Finally, we have
$\sum_{j=1}^{\ell}h(|I_{j}|)\geq\sum_{j=1}^{\ell}h(\varepsilon_{v(j)+1})\geq\sum_{r\geq r_{0}}v_{r}\cdot h(\varepsilon_{r+1}),$
where $r_{0}=\min\{r\mid k_{r}>0\}$. Since $\sum_{r\geq r_{0}}(m_{r}/N_{r})\geq 1/2$, it follows that
$\sum_{j=1}^{\ell}h(|I_{j}|)\geq\sum_{r\geq r_{0}}\frac{m_{r}}{N_{r}}\cdot N_{r}\cdot h(\varepsilon_{r+1})\geq\frac{1}{2}\cdot\min_{r\geq r_{0}}(N_{r}\cdot h(\varepsilon_{r+1})),$
which tends to $\infty$ when $r_{0}\to+\infty$. But when the norm of the covering tends to 0, we do have $r_{0}\to+\infty$, and the proof is complete. ∎
We now proceed to prove Theorem 1.1. The proof will be divided into the following three lemmas.
Lemma 3.4.
Let $h$ be a dimension function. If
$\lim_{\varepsilon\to\infty}\frac{\log h(\varepsilon)}{\log\varepsilon}=0$
then $\mathcal{H}^{h}(k^{-1}(3))=\infty$.
Proof.
The goal is to show that $\mathcal{H}^{h}(k^{-1}(3))=\infty$ under the given assumptions on the function $h$. To achieve this, we construct a Cantor subset $C\subset k^{-1}(3)$ such that $\mathcal{H}^{h}(C)=\infty$. The construction will rely on a recursive sequence that reflects the asymptotic behavior of $h$ near zero.
Let $c:=\log(3+2\sqrt{2})>1$. We now define the sequences $\{f_{i}\}$, $\{\epsilon_{i}\}$, and $\{\delta_{i}\}$. We begin with the initial values:
$f_{1}=3,\qquad\epsilon_{1}=e^{-cf_{1}}.$
We define $\delta_{2}$ as follows:
$\delta_{2}=\sup_{0\leq\epsilon\leq\epsilon_{1}}\frac{|\log(h(\epsilon))|}{|\log(\epsilon)|}.$
Given $\delta_{k}$, we define $f_{k}$, then $\epsilon_{k}$, and finally $\delta_{k+1}$ by the following recursion for $k\geq 2$:
$f_{k}=\left\lfloor\frac{1}{2c\delta_{k}}\right\rfloor,\qquad\epsilon_{k}=e^{-c\sum_{i=1}^{k}f_{i}},\qquad\delta_{k+1}=\sup_{0\leq\epsilon\leq\epsilon_{k}}\frac{|\log(h(\epsilon))|}{|\log(\epsilon)|},$
where $\lfloor\cdot\rfloor$ denote the floor function. Note that, since $\lim_{\epsilon\to 0}\frac{|\log(h(\epsilon))|}{|\log(\epsilon)|}=0$ we have that $\delta_{k}\to 0$ and $f_{k}\to\infty$ as $k\to\infty$. Now we construct a Cantor set determined by the sequences defined above. For each $k$, let $r_{i}\in\mathbb{N}$ for $i=1,\dots,k$ satisfy
$f_{i}\leq r_{i}<2f_{i}.$
We consider the set of real numbers whose continued fraction expansion begins with the following pattern:
[…]
That is, the expansion consists of alternating blocks of $r_{i}$ ones followed by two 2’s, for $i=1$ to $k$. This defines a finite-level construction of a Cantor set based on the growth of the sequences $(f_{i})$. Let $C$ be the Cantor set obtained by repeating the above construction for all $n\in\mathbb{N}$. Then the set $C$ satisfies the following properties:
  1. (1)
    $C\subset k^{-1}(3)$. This follows from the fact that the sequence $(r_{i})$ is increasing. Indeed, since by construction $2f_{i}=2\left\lfloor\frac{1}{2c\delta_{i}}\right\rfloor\leq\left\lfloor\frac{2}{2c\delta_{i}}\right\rfloor\leq f_{i+1}$, we have that $r_{i}\leq 2f_{i}\leq f_{i+1}\leq r_{i+1}$, so the number of 1’s between each pair of 2’s increases. Therefore, any $x\in C$ is attainable and $k(x)=3$ because of Lemma 3.2.
  2. (2)
    At step $n$ of the construction, the set $C$ consists of exactly $N_{n}:=\prod_{i=1}^{n}f_{i}$ closed intervals. This is because for each index $i$, the repetition number $r_{i}$ ranges over exactly $f_{i}$ possible values, independent from the other levels.
  3. (3)
    Each closed interval appearing in step $n$ of the construction has length at least $\epsilon_{n}$. This is a classical fact in the theory of continued fractions (see (4)).
Now, since $f_{i}\leq\frac{1}{2c\delta_{i}}$, we have that $2c\,\delta_{i}\,f_{i}\leq\log(f_{i})$. Therefore, we have:
$\log(N_{n}h(\epsilon_{n+1}))$ $=\log(N_{n})+\log(h(\epsilon_{n+1}))$
$\geq\sum_{i=1}^{n}\log(f_{i})-c\delta_{n+2}\sum_{i=1}^{n+1}f_{i}$
$\geq\left(1-\frac{1}{2}\right)\sum_{i=1}^{n}\log f_{i}-c\delta_{n+2}f_{n+1}$
$\geq\frac{1}{2}\sum_{i=1}^{n}\log f_{i}-c$
This last expression tends to infinity as $n\to\infty$, which implies that $N_{n}h(\epsilon_{n+1})\to\infty$. Consequently, we have:
$\mathcal{H}^{h}(C)=\mathcal{H}^{h}(k^{-1}(3))=\infty$
by using Lemma 3.3. ∎
Lemma 3.5.
If $\lim_{\varepsilon\to 0}\frac{\log h(\varepsilon)}{\log\varepsilon}=0$, then $k^{-1}(3)$ does not have $\sigma$-finite $\mathcal{H}^{h}$-measure.
Proof.
Define the dimension function $g(x)=h(x)^{2}$. We have that
$\lim_{\varepsilon\to 0}\frac{\log g(\varepsilon)}{\log\varepsilon}=\lim_{\varepsilon\to 0}\frac{2\log h(\varepsilon)}{\log\varepsilon}=0.$
Hence $\mathcal{H}^{g}(k^{-1}(3))=\infty$. Moreover
$\lim_{\varepsilon\to 0}\frac{g(\varepsilon)}{h(\varepsilon)}=\lim_{\varepsilon\to 0}h(\varepsilon)=0.$
In particular, if $k^{-1}(3)=\cup_{n}E_{n}$ with $\mathcal{H}^{h}(E_{n})<\infty$, we must have $\mathcal{H}^{g}(E_{n})=0$. But this would imply that
$\mathcal{H}^{g}(k^{-1}(3))=\mathcal{H}^{g}(\cup_{n}E_{n})\leq\sum_{n}\mathcal{H}^{g}(E_{n})=0,$
which is clearly a contradiction. ∎
Lemma 3.6.
Let $h$ be a dimension function. If
$\limsup_{\varepsilon\to\infty}\frac{\log h(\varepsilon)}{\log\varepsilon}>0$
then $\mathcal{H}^{h}(k^{-1}(3))=0$.
Proof.
By hypothesis, there is a positive $d$ and a decreasing sequence $\varepsilon_{k}$ such that $\varepsilon_{k}\to 0$ when $k\to\infty$ such that
(8) $h(\varepsilon_{k})\leq\varepsilon_{k}^{d},\quad\text{for all $k$}.$
Take $m\in\mathbb{N}$ so large so that $9m^{3}<2^{dm-1}$.
Define the sequence of positive integers $r_{k}=m\left\lceil\frac{|\log_{2}\varepsilon_{k}|}{m}\right\rceil$ for all $k\geq 1$.
By Lemma 2.4, we know there is $n_{0}=n_{0}(x)\geq 2$ and a sequence of positive integers $(b_{j})_{j<n_{0}}$ such that
(9) $(\dots,b_{n_{0}-2},b_{n_{0}-1},a_{n_{0}},a_{n_{0}+1},a_{n_{0}+2},\dots)\in\Sigma_{3+6^{-m}}.$
In particular, this shows that $k^{-1}(3)$ can be covered by the countable union
$k^{-1}(3)\subset\bigcup_{n_{0}\geq 2}\bigcup_{(a_{1},\dots,a_{n_{0}-1})\in(\mathbb{N}^{*})^{n_{0}-1}}\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}.$
where $K_{t}$ was defined on (1). If we show that $\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}$ has zero Hausdorff hh–measure, then $\mathcal{H}^{h}(k^{-1}(3))=0$. From now on we fix $(a_{1},\dots,a_{n_{0}-1})$ and assume that $x\in\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}$.
It follows rom (9) that given any sub-block of length $m$ at a position $n\geq n_{0}$ of $x=[0;a_{1},a_{2},\dots]$, that is, a word of the form $(a_{n},\dots,a_{n+m-1})$, it must belong to $\Sigma(3,m)$ and we have at most $9m^{3}<2^{md-1}$ possibilities for such sub-blocks.
We want to build a coverings of $k^{-1}(3)$ with arbitrarily small hh–measure. For this, we will employ Theorem 2.1 to cover each $\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}$.
Observe that we can divide the block $w=(a_{n_{0}},a_{n_{0}+1},\dots,a_{n_{0}+r_{k}-1})$ in the subblocks $(a_{n_{0}+im},\dots,a_{n_{0}+(i+1)m-1})$ for $i=0,\dots,r_{k}/m-1$. In particular, we can cover the block $B$ with at most
$(2^{md-1})^{r_{k}/m}$ $=(2^{md-1})^{\lceil|\log_{2}\varepsilon_{k}|/m\rceil}$
$\leq(2^{md-1})^{|\log_{2}\varepsilon_{k}|/m+1}$
$=(2^{|\log_{2}\varepsilon_{k}|})^{d}\cdot 2^{-|\log_{2}\varepsilon_{k}|/m}\cdot 2^{md-1}$
$=\varepsilon_{k}^{-d}\cdot 2^{-|\log_{2}\varepsilon_{k}|/m}\cdot 2^{md-1}$
intervals $I(p,w)$ where $p=(a_{1},\dots,a_{n_{0}-1})$ is fixed and $w\in\Sigma(3,n)^{r_{k}/m}$. The length of the intervals $I(p,w)$ is at most $2^{-n_{0}-r_{k}}\leq 2^{-r_{k}}\leq 2^{-|\log_{2}\varepsilon_{k}|}=\varepsilon_{k}$. Using (8), we have that the measure of the covering $I(p,w)$ is at most
$\sum_{w\in\Sigma(3,n)^{r_{k}/m}}h(\operatorname{diam}I(w))$ $\leq\sum_{w\in\Sigma(3,n)^{r_{k}/m}}h(\varepsilon_{k})$
$\leq(2^{md-1})^{r_{k}/m}\cdot\varepsilon_{k}^{d}$
$\leq\varepsilon_{k}^{-d}\cdot 2^{-|\log_{2}\varepsilon_{k}|/m}\cdot 2^{md-1}\cdot\varepsilon_{k}^{d}$
$=2^{md-1}\cdot 2^{-|\log_{2}\varepsilon_{k}|/m}.$
Since the sequence $\varepsilon_{k}$ is decreasing and $2^{md-1}$ is constant, the above term goes to zero as $k\to\infty$, which shows that $\mathcal{H}^{h}(k^{-1}(3))=0$.
We would like to remark that our result can be compared with those obtained by Olsen and Renfro in [8, 9], concerning the Hausdorff measure of the set of Liouville numbers $\mathbb{L}$. In these articles, the authors proved the following
Theorem 3.7.
Let $h$ be a dimension function. Then:
  1. (1)
    If $\limsup_{r\to 0}\frac{\Gamma_{h}(r)}{r^{t}}=0$ for some $t>0$, then $\mathcal{H}^{h}(\mathbb{L})=0$.
  2. (2)
    If $\limsup_{r\to 0}\frac{\Gamma_{h}(r)}{r^{t}}>0$ for all $t>0$, then the set $\mathbb{L}$ does not have $\sigma$-finite $\mathcal{H}^{h}$-measure.
Here, $\Gamma_{h}(r)=r\cdot\inf_{0<s\leq r}\frac{h(s)}{s}$ and $\mathbb{L}$ denotes the set of Liouville numbers.
It is proved in [9, Lemma 2.2] that for all dimension functions $h$ and all sets $E\subset\mathbb{R}$, we have:
$\mathcal{H}^{\Gamma_{h}}(E)\leq\mathcal{H}^{h}(E)\leq 2\mathcal{H}^{\Gamma_{h}}(E).$
Using this, we may reformulate Theorem 1.1 as follows,
Theorem 3.8.
Let $h$ be a dimension function. Then:
  1. (1)
    If $\liminf_{r\to 0}\frac{\Gamma_{h}(r)}{r^{t}}=0$ for some $t>0$, then $\mathcal{H}^{h}(k^{-1}(3))=0$.
  2. (2)
    If $\liminf_{r\to 0}\frac{\Gamma_{h}(r)}{r^{t}}>0$ for all $t>0$, then the set $k^{-1}(3)$ does not have $\sigma$-finite $\mathcal{H}^{h}$-measure.
This comparison shows the set of Liouville numbers is strictly larger than $k^{-1}(3)$ from the point of view of Hausdorff dimension, since there are dimension functions $h$ for which $\mathcal{H}^{h}(\mathbb{L})=\infty$ while $\mathcal{H}^{h}(k^{-1}(3))=0$. Indeed, consider for example the right continuous dimension function
$h(\varepsilon)=\begin{cases}e^{-(2k+1)!}\cdot\left(e^{(2k)!}\right)^{1-1/k},&\text{if }e^{-(2k+2)!}\leq\varepsilon\leq e^{-(2k+1)!};\\ \varepsilon\cdot\left(e^{(2k)!}\right)^{1-1/k},&\text{if }e^{-(2k+1)!}\leq\varepsilon<e^{-(2k)!}.\end{cases}$
It is easy to see that
$\Gamma_{h}(\varepsilon)=\begin{cases}\varepsilon\cdot\left(e^{(2k)!}\right)^{1-1/k},&\text{if }e^{-(2k+1)!}\leq\varepsilon\leq\left(e^{(2k)!}\right)^{-1+1/k}\cdot h(e^{-(2k-1)!});\\ h(e^{-(2k-1)!}),&\text{if }\left(e^{(2k)!}\right)^{-1+1/k}\cdot h(e^{-(2k-1)!})\leq\varepsilon<e^{-(2k-1)!}.\end{cases}$
In particular, for $\varepsilon_{k}=e^{-(2k+1)!}$ we have that
$\lim_{k\to\infty}\frac{\log h(\varepsilon_{k})}{\log\varepsilon_{k}}=1-\lim_{k\to\infty}\frac{(2k)!(1-1/k)}{(2k+1)!}=1,$
which implies $\limsup_{\varepsilon\to 0}\frac{\log h(\varepsilon)}{\log\varepsilon}>0$ whence $\mathcal{H}^{h}(k^{-1}(3))=0$ because of Theorem 1.1. On the other hand, for $\varepsilon_{k}=e^{-(2k)!-1}$ and any $t>0$ one has
$\lim_{k\to\infty}\frac{\Gamma_{h}(\varepsilon_{k})}{\varepsilon_{k}^{t}}=\lim_{k\to\infty}e^{t-1}\left(e^{(2k)!}\right)^{1-1/k+t-1}=\lim_{k\to\infty}e^{t-1}\left(e^{(2k)!}\right)^{t-1/k}=\infty,$
which implies $\limsup_{\varepsilon\to 0}\frac{\Gamma_{h}(\varepsilon)}{\varepsilon^{t}}>0$ for all $t>0$, so $\mathcal{H}^{h}(\mathbb{L})=\infty$ because of Theorem 3.7.
Finally, let us mention that from the topological point of view, the Liouville numbers are also larger than $k^{-1}(3)$. First, we will show that $k^{-1}(3)$ is not $F_{\sigma}$. Indeed, given some knowledge about the generalized Hausdorff measures of a set implies some restriction on its topology, so the following lemma can be of independent interest.
Lemma 3.9.
Suppose that $E\subset\mathbb{R}$ is a bounded set such that for any dimension function $h$ with
(10) $\lim_{\varepsilon\to 0}\frac{\log h(\varepsilon)}{\log\varepsilon}=0,$
one has $\mathcal{H}^{h}(E)>0$. Then $E$ is not $F_{\sigma}$.
Proof.
Suppose by contradiction that $E=\bigcup_{n}F_{n}$ for some compact sets $F_{n}$. For each positive integer $n$, we will construct now a dimension function $h_{n}$ such that $\mathcal{H}^{h_{n}}(F_{n})=0$ and
(11) $\lim_{\varepsilon\to 0}\frac{\log h_{n}(\varepsilon)}{\log\varepsilon}=0.$
Fix $n$ and denote $F=F_{n}$, where we dropped the dependence on $n$ to simplify the following notation. Since $F$ is a compact set with Hausdorff dimension zero, by definition, there is a finite open covering $\{U_{1,i}\}_{i}$ of $F$ such that $\sum_{i}|U_{1,i}|<1$. Let $r_{1}=\min_{i}|U_{1,i}|$ and $s_{1}=\max_{i}|U_{1,i}|$. Assuming $r_{1},\dots,r_{k},s_{1},\dots,s_{k}$ are defined, by the same reason above, there is a finite open covering $\{U_{k+1,i}\}_{i}$ of $F$ such that
(12) $\sum_{i}|U_{k+1,i}|^{1/(k+1)}<r_{k},\qquad\text{and}\qquad\max_{i}|U_{k+1,i}|<r_{k}^{1+1/k},$
so define $r_{k+1}=\min_{i}|U_{k+1,i}|$ and $s_{k+1}=\max_{i}|U_{k+1,i}|$. Finally, let $h_{n}$ be the right continuous non-decreasing function
$h_{n}(\varepsilon)=\begin{cases}\varepsilon^{1/k},&\text{if }r_{k}\leq\varepsilon<s_{k};\\ r_{k}^{1/k},&\text{if }s_{k+1}\leq\varepsilon\leq r_{k}.\end{cases}$
Note that $r_{1}\leq s_{1}<1$ since $\sum_{i}|U_{1,i}|<1$. Since (12) implies $r_{k+1}\leq s_{k+1}<r_{k}^{1+1/k}$ for all $k\geq 1$, it follows by induction that $r_{k}\leq s_{k}<1$ for all $k\geq 1$ and also $r_{k+1}\leq r_{1}^{(1+1/2)\dotsb(1+1/k)}=r_{1}^{(k+1)/2}$. Since $r_{1}<1$, this shows that $r_{k},s_{k}\to 0$ as $k\to\infty$, so indeed $h_{n}$ is a non-decreasing function such that $\lim_{\varepsilon\to 0}h_{n}(\varepsilon)=0$. The fact that $\mathcal{H}^{h_{n}}(F)=0$ is a consequence of the construction: given any $\delta>0$, we let $k$ be sufficiently large so that $r_{k}^{1+1/k}<\delta$, so using (12) with the covering $\{U_{k+1,i}\}_{i}$ of $F$ gives that
$\sum_{i}h(\operatorname{diam}U_{k+1,i})=\sum_{i}|U_{k+1,i}|^{1/(k+1)}<r_{k}.$
Since $r_{k}\to 0$ as $k\to\infty$, we conclude that $\mathcal{H}^{h_{n}}(F)=0$. Now to prove (11), just observe that for $s_{k+1}\leq\varepsilon<s_{k}$ one has that $0<\log h_{n}(\varepsilon)/\log\varepsilon\leq 1/k$, so (11) follows from the fact that $s_{k}\to 0$ as $k\to\infty$.
Now that we have constructed dimension functions $h_{n}$ that satisfy (11) and $\mathcal{H}^{h_{n}}(F_{n})=0$, we will build another dimension function $h$ such that for each $n$, we have $h(\varepsilon)\leq h_{n}(\varepsilon)$ for all $\varepsilon>0$ sufficiently small but
$\lim_{\varepsilon\to 0}\frac{\log h(\varepsilon)}{\log\varepsilon}=0.$
In particular, it follows that $\mathcal{H}^{h}(F_{n})\leq\mathcal{H}^{h_{n}}(F_{n})=0$ for all $n$, but this is a contradiction to the hypothesis (10) since $\mathcal{H}^{h}(E)\leq\sum_{n}\mathcal{H}^{h}(F_{n})=0$.
To construct such a function, given any positive integer $n$, let $0<\delta_{n}<\delta_{n-1}$ be such that
$\sup_{0<\varepsilon\leq\delta_{n}}\left\{\frac{\log h_{1}(\varepsilon)}{\log\varepsilon},\dots,\frac{\log h_{n}(\varepsilon)}{\log\varepsilon}\right\}<\frac{1}{n}.$
Finally, define the right continuous dimension function $h(\varepsilon)=\min\{h_{1}(\varepsilon),\dots,h_{n}(\varepsilon)\}$ for $\delta_{n+1}\leq\varepsilon<\delta_{n}$. It is easy to check that $h$ satisfies the claimed properties, so the proof that $E$ is not $F_{\sigma}$ is done.
Lemma 3.10.
The set $k^{-1}(3)$ is not $G_{\delta}$ and not $F_{\sigma}$, but it is $F_{\sigma\delta}$.
Proof.
In general, for $t\in[0,\infty]$ such that $k^{-1}(t)$ is non-empty, the set $k^{-1}(t)$ is not a $G_{\delta}$ set, i.e., is not the intersection of countably many open sets. The reason why $k^{-1}(t)$ is not $G_{\delta}$ is because for any $s\in\mathbb{R}$, we have that
$k^{-1}[s,\infty]=\bigcap_{n\geq 1}\bigcup_{q\geq 2}\bigcup_{p\in\mathbb{Z}}\left(\frac{p}{q}-\frac{1}{(s-1/n)q^{2}},\frac{p}{q}+\frac{1}{(s-1/n)q^{2}}\right)$
is a dense $G_{\delta}$ set. If $k^{-1}(t)$ is $G_{\delta}$, then $k^{-1}(t)\cap k^{-1}[t+1,\infty]=\varnothing$ would be a dense $G_{\delta}$ set, being the intersection of two $G_{\delta}$ dense sets, a contradiction.
Now we prove that the set $k^{-1}(3)$ is an $F_{\sigma\delta}$ set: it is the countable intersection of $F_{\sigma}$ sets where each $F_{\sigma}$ is the countable union of closed sets with empty interior. Indeed, since $k^{-1}[3+1/m,\infty]$ is a $G_{\delta}$ set, we have that $k^{-1}(-\infty,3+1/m)=\bigcup_{n\geq 1}X_{m,n}$ where each $X_{m,n}$ is a closed set with empty interior (since $k^{-1}(-\infty,t)$ contains no rational number for any $t\in\mathbb{R}$). On the other hand, the set $k^{-1}(-\infty,3)$ is actually countable, because $\{t<3:k^{-1}(t)\neq\varnothing\}=\{\ell_{1}=\sqrt{5}<\ell_{2}=2\sqrt{2}<\dots\}$ is countable and each $k^{-1}(\ell_{n})$ is also countable because of Markov’s theorem [2, Theorem 16]. In particular, since the complement of a point is the countable union of closed sets, the complement of the countable set $k^{-1}(-\infty,3)$ is an $F_{\sigma\delta}$, that is, it can be written as $k^{-1}[3,\infty]=\mathbb{R}\setminus k^{-1}(-\infty,3)=\bigcap_{m\geq 1}\bigcup_{n\geq 1}Y_{m,n}$ for some closed sets $Y_{m,n}$. Finally, we have that
$k^{-1}(3)=k^{-1}(-\infty,3]\cap k^{-1}[3,\infty)=\bigcap_{m\geq 1}\bigcup_{n_{1},n_{2}\geq 1}(X_{m,n_{1}}\cap Y_{m,n_{2}})$
where each $X_{m,n_{1}}\cap Y_{m,n_{2}}$ is a closed set with empty interior.
Finally, notice that if $k^{-1}(3)$ is $F_{\sigma}$ then $k^{-1}(3)\cap[0,1]$ would be a bounded $F_{\sigma}$ set, but by Lemma 3.4 we have $\mathcal{H}^{h}(k^{-1}(3)\cap[0,1])=\infty$ (since $k^{-1}(3)$ is invariant by integer translations), which is a contradiction to Lemma 3.9.
The previous lemma contrasts with the situation for Liouville numbers, which form a dense $G_{\delta}$ set. Another major difference is that the set of Liouville numbers is invariant under translation by rational numbers. It is unsurprising that $k^{-1}(3)$ is not invariant by such translations, since adding a rational number to a continued fraction can drastically alter its coefficients. To illustrate this with a concrete example, let $e_{m}=6\cdot\lceil m/12\rceil$ and consider the continued fraction
$x=[0;1^{e_{1}},2,2,1^{e_{2}},2,2,1^{e_{3}},2,2,\dots].$
By Lemma 3.2 we have that $x\in k^{-1}(3)$. If we denote by $p_{n}/q_{n}$ the continued fraction of $x$, then it is an elementary exercise to show that for all $m\geq 1$
$q_{2+e_{1}+\dots+e_{6m}+8m}\equiv 2\pmod{4}\qquad\text{and}\qquad p_{2+e_{1}+\dots+e_{6m}+8m}\equiv 1\pmod{4}$.
In particular, for $n=2+e_{1}+\dots+e_{6m}+8m$, letting
$y=x+1/2,\qquad\text{and}\qquad(p,q)=((p_{n}-q_{n}/2)/2,q_{n}/2),$
we will have that the inequality
$\left|y-\frac{p}{q}\right|=\left|\left(x+\frac{1}{2}\right)-\left(\frac{p_{n}}{q_{n}}-\frac{1}{2}\right)\right|=\left|x-\frac{p_{n}}{q_{n}}\right|<\frac{1}{q_{n}^{2}}=\frac{1}{4q^{2}},$
has infinitely many solutions $p/q\in\mathbb{Q}$. By the definition of the function $k$, we have that $k(x+1/2)=k(y)\geq 4$ (in fact $k(y)\geq 12$).
4. The set $K_{3}$
Lemma 4.1.
Let $h$ be a dimension function. If
$\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>3$
then $\mathcal{H}^{h}(K_{3})=0$.
Proof.
By hypothesis, there is a positive $d>0$ and a decreasing sequence $\varepsilon_{k}$ such that $\varepsilon_{k}\to 0$ when $k\to\infty$ such that
(13) $h(\varepsilon_{k})\leq\frac{1}{|\log\varepsilon_{k}|^{3+d}},\quad\text{for all $k$}.$
Given $k$, let $n_{k}=\lceil\log_{2}\varepsilon_{k}\rfloor$. Note that for any $n_{k}\geq 1$, the set $\{I(w):w=(a_{1},\dots,a_{n_{k}})\in\Sigma(3,n_{k})\}$ is a covering of $K_{3}$ by closed intervals. Since $\operatorname{diam}I(w)<2^{-n_{k}}$, we have that
$\sum_{w\in\Sigma(3,n_{k})}h(\operatorname{diam}I(w))$ $\leq|\Sigma(3,n_{k})|h(\varepsilon_{k})$
$\leq\frac{9n_{k}^{3}}{|\log\varepsilon_{k}|^{3+d}}\leq\frac{9n_{k}^{3}}{(\log(2)-1)^{3+d}n_{k}^{3+d}}=\frac{9}{(\log(2)-1)^{3+d}n_{k}^{d}}.$
Since $n_{k}\to\infty$ as $k\to\infty$, this proves the lemma. ∎
Acknowledgements
Nicolas Angelini gratefully acknowledges the Instituto de Matemática Pura e Aplicada (IMPA) for hosting his research visit, and the Academia de Ciencias de América Latina (ACAL) for financial support during this stay.
Funding
The research of the first author was partially supported by CNPq and FAPERJ. The resarch of the second author was partially supported by CAPES and FAPERJ. The research of the third author was supported by Programa de Subvenciones para Pasantías Cortas de Investigadores Jóvenes of the Academia de Ciencias de América Latina (ACAL), and was also partially supported by grants PICT 2022-4875 (ANPCyT), PIP 202287/22 (CONICET), and PROICO 3-0720 “Análisis Real y Funcional. Ec. Diferenciales”.

### v2

@@
 $\mathcal{H}^{h}(E)=\lim_{\delta\to 0}\inf\left\{\sum_{i}h(\operatorname{diam}(U_{i})):E\subset\bigcup_{i}U_{i},\,\operatorname{diam}(U_{i})<\delta\right\},$
-where $h:(0,\infty)\to(0,\infty)$ is an nondecreasing and right continuous function such that $h(r)\to 0$ as $r\to 0$.
+where $h:(0,\infty)\to(0,\infty)$ is a nondecreasing and right continuous function such that $h(r)\to 0$ as $r\to 0$.
 Such a function is called a dimension function or gauge function.
@@
 The image $\{k(x):x\in\mathbb{R}\setminus\mathbb{Q},k(x)<\infty\}$ of this function is known as the Lagrange spectrum.
-It is well known that $k^{-1}(3)$ is an uncountable dense set and it is a consequence of [7, Theorem 1] that $k^{-1}(3)$ has Hausdorff dimension 0. Indeed, for $t\in\mathbb{R}$ define
-(1) $K_{t}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}^{*})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq t,\forall k\in\mathbb{Z}\}.$
+It is well known that $k^{-1}(3)$ is an uncountable dense set and it is a consequence of [10, Theorem 1] that $k^{-1}(3)$ has Hausdorff dimension 0. Indeed, for $t\in\mathbb{R}$ define
+(1) $K_{t}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}_{>0})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq t,\forall k\in\mathbb{Z}\}.$
 Then $K_{t}\subset k^{-1}(-\infty,t]$ and $D(t)=\dim_{H}(K_{t})=\overline{\dim_{B}}(K_{t})=\dim_{H}(k^{-1}(-\infty,t])$ is a continuous function with $\max\{t\in\mathbb{R}:D(t)=0\}=3$.
@@
 We will provide an explicit example of a dimension function $h$ such that $\mathcal{H}^{h}(\mathbb{L})=\infty$ while $\mathcal{H}^{h}(k^{-1}(3))=0$.
-A more general property of Liouville numbers is that they are immeasurable, that is, any translation invariant Borel measure on $\mathbb{R}$ either assigns zero measure to $\mathbb{L}$ or is not $\sigma$–finite on $\mathbb{L}$ [3, Theorem 1.1] (note that this includes the Hausdorff hh–measures).
+A more general property of Liouville numbers is that they are immeasurable, that is, any translation invariant Borel measure on $\mathbb{R}$ either assigns zero measure to $\mathbb{L}$ or is not $\sigma$–finite on $\mathbb{L}$ [4, Theorem 1.1] (note that this includes the Hausdorff hh–measures).
 The main result of [4] applies to a broad class of sets, however, it does not apply to $k^{-1}(3)$ because of several reasons.
@@
     If
-    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>3$
+    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}<3,$
+    then $\mathcal{H}^{h}(K_{3})=\infty$.
+  2. (2)
+    If
+    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>3,$
     then $\mathcal{H}^{h}(K_{3})=0$.
-A forthcoming work of the first two authors with Zhe Cao will imply the following result: for any $x=[x_{0};x_{1},x_{2},\dots]\in k^{-1}(3)$ such that the inequality $\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}$ has only finitely many solutions, there is an $N\in\mathbb{N}$ such that $0;x_{N},x_{N+1},\dots]\in K_{3}$.
-In particular, if we denote by $B_{3}$ the set of non–attainable elements of $k^{-1}(3)$, then for example for the dimension function $h(x)=\frac{1}{|\log\varepsilon|^{4}}$ one has that $\mathcal{H}^{h}(k^{-1}(3))=\infty$ but $\mathcal{H}^{h}(B_{3})=0$.
-In conclusion, there is a large gap between the generalized Hausdorff dimension of these sets, so the set of non–attainable elements $x\in k^{-1}(3)$ is also significantly smaller than $k^{-1}(3)$.
+In fact the previous theorem is corollary of a stronger result.
+Theorem 1.5.
+For the dimension function
+$g(\varepsilon)=\frac{\log|\log\varepsilon|}{|\log\varepsilon|^{3}(\log\log|\log\varepsilon|)},$
+we have that $\mathcal{H}^{g}(K_{3})<\infty$.
+For a large enough constant $\hat{C}>0$, the dimension function
+$g(\varepsilon)=\frac{\exp(\hat{C}\sqrt{\log|\log\varepsilon|})}{|\log\varepsilon|^{3}}$
+satisfies $\mathcal{H}^{g}(K_{3})=\infty$.
+Then for example for the dimension function $h(\varepsilon)=\frac{1}{|\log\varepsilon|^{3}}$ one has that $\mathcal{H}^{h}(k^{-1}(3))=\infty$ but $\mathcal{H}^{h}(K_{3})=0$.
+In conclusion, there is a large gap between the generalized Hausdorff dimension of these sets, so the set $K_{3}$ is significantly smaller than $k^{-1}(3)$.
+Denote by $B_{3}$ the set of non–attainable elements of $k^{-1}(3)$, that is
+$B_{3}=\left\{x\in\mathbb{R}\setminus\mathbb{Q}:\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}\text{ has only finitely many solutions }\frac{p}{q}\in\mathbb{Q}\right\}.$
+A recent work [3] by the first two authors with Zhe Cao imply the following result: for any $x=[x_{0};x_{1},x_{2},\dots]\in B_{3}$, there is an $N\in\mathbb{N}$ such that $0;x_{N+1},x_{N+2},\dots]\in K_{3}$.
+However, while the Cantor set $K_{3}$ is forward invariant by the Gauss map, the set $B_{3}$ is not.
+The next theorem shows that the non–attainable elements $B_{3}$ are even smaller than $K_{3}$ and thus smaller than $k^{-1}(3)$.
+Theorem 1.6.
+Let $h$ be a dimension function.
+  1. (1)
+    If
+    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}<2,$
+    then $\mathcal{H}^{h}(B_{3})=\infty$.
+  2. (2)
+    If
+    $\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>2,$
+    then $\mathcal{H}^{h}(B_{3})=0$.
+The proof of the previous theorem is essentially the same as Theorem 1.4.
+The key distinction is that, in the case of $K_{3}$, one must account for permutations of renormalizable words, reflecting its invariance under the Gauss map.
+By contrast, the structure of $B_{3}$ is considerably more rigid so this additional counting is not required.
+For completeness, one could also formulate an analogue of Theorem 1.5 for $B_{3}$, but we refrain from doing so here in order to avoid unnecessary repetition.
+From this version it follows that $\mathcal{H}^{h}(B_{3})=0$ for the gauge function $h(\varepsilon)=\frac{1}{|\log\varepsilon|^{2}}$.
 2. Preliminaries
+2.1. Continued fraction and Hausdorff dimension
 Given an nondecreasing and right continuous function $h:[0,\infty)\to[0,\infty)$ with $h(0)=0$, we define the corresponding Hausdorff measure $\mathcal{H}^{h}$ as
@@
 We say that a set $E$ is an hh–set if $0<\mathcal{H}^{h}(E)<\infty$. It is not always possible to find a dimension function for which this holds.
-Given $\underline{a}\in(\mathbb{N}^{*})^{\mathbb{Z}}$ define
+Given $\underline{a}\in(\mathbb{N}_{>0})^{\mathbb{Z}}$ define
 $\lambda_{i}((a_{n})_{n\in\mathbb{Z}})=[a_{i};a_{i+1},a_{i+2},\dots]+[0;a_{i-1},a_{i-2},\dots].$
@@
 (2) $\left|x-\frac{p_{n}}{q_{n}}\right|=\frac{1}{(\gamma_{n+1}+\eta_{n+1})q_{n}^{2}},$
-where $\gamma_{n+1}=[a_{n+1};a_{n+2},\dots]$ and $\eta_{n+1}=[0;a_{n},a_{n-1},\dots,a_{1}]$. In particular we have the formula $k(x)=\limsup_{i\to\infty}\lambda_{i}(x)$.
+where $\gamma_{n+1}=[a_{n+1};a_{n+2},\dots]$ and $\eta_{n+1}=[0;a_{n},a_{n-1},\dots,a_{1}]$. In particular we have the formula $k(x)=\limsup_{n\to\infty}\lambda_{n}(x)$.
 A curious and crucial identity to study irrational numbers $x$ with values $k(x)$ near 3 has been the following:
 (3) $2;2,z]+[0;1,1,z]=3,\qquad\text{for all $z>0$}.$
-Given a finite word of positive integers $w=(a_{1},\dots,a_{n})\in(\mathbb{N}^{*})^{n}$, define the closed sub-interval of $[0,1]$
+Given a finite word of positive integers $w=(a_{1},\dots,a_{n})\in(\mathbb{N}_{>0})^{n}$, define the closed sub-interval of $[0,1]$
 $I(w):=\{x\in[0,1]\ \mid\ x=[0;a_{1},a_{2},\dots,a_{n},x],x\geq 1\}\cup\{[0,a_{1},a_{2},\dotsc,a_{n}]\},$
@@
 (4) $(3+2\sqrt{2})^{-n-1}<\operatorname{diam}I(w)<\left(\frac{3+\sqrt{5}}{2}\right)^{-n+1}$
+The well known estimate of Jarnik [7] says that the Hausdorff dimension of the Gauss-Cantor set $C(n)=\{[0;a_{1},a_{2},\dots]:1\leq a_{i}\leq n,\forall i\geq 1\}$ satisfies $1-C_{2}/n<\dim_{H}(C(n))<1-C_{1}/n$ for some explicit constants $0<C_{1}<C_{2}$ and all $n$ large.
+For our purposes, Jarnik result is enough, however a much more precise asymptotic estimate is due to Hensley [6]:
+(5) $\dim_{H}(C(n))=1-\frac{6}{\pi^{2}n}-\frac{72\log n}{\pi^{4}n^{2}}+O\left(\frac{1}{n^{2}}\right).$
+Another related result by Hensley estimates the cardinality of the set
+[…]
+Hensley [6] proved that for any $n>1$, there is a constant $E_{n}>0$ such that $\lim_{x\to\infty}x^{-2\dim_{H}(C(n))}|Q_{n}(x)|=E_{n}$.
+For our purposes, we will need a more elementary estimate that is not dependent on the rate of convergence.
+Proposition 2.1.
+For all $x>1$,
+(6) $x^{2\dim_{H}(C(n))}\leq Q_{n}(x)\leq 2x^{2\dim_{H}(C(n))}.$
+Proof.
+Recall that a Gauss-Cantor set $K(B)$ is defined by an alphabet $B=\{\beta_{1},\dots,\beta_{\ell}\}$ and for each word $\beta_{j}\in(\mathbb{N}_{>0})^{r_{j}}$, the interval $I_{j}=I(\beta_{j})$ and $\psi|_{I_{j}}:=G^{r_{j}}|_{I_{j}}$ is an iterate of the Gauss map $G(x)=\frac{1}{x}-\lfloor\frac{1}{x}\rfloor$.
+This defines an expanding map $\psi\colon I_{1}\cup\dotsb\cup I_{\ell}\to I$ with Markov partition $I_{1},\dots,I_{\ell}$.
+According to Palis–Takens [13, Chapter 4, Pages 68-71], if we let
+$\lambda_{j}=\inf|\psi^{\prime}|_{I_{j}}|,\qquad\Lambda_{j}=\sup|\psi^{\prime}|_{I_{j}}|$
+and $d_{1},d_{2}\geq 0$ be such that
+$\sum_{i=1}^{\ell}\lambda_{j}^{-d_{2}}=1,\qquad\sum_{i=1}^{\ell}\Lambda_{j}^{-d_{1}}=1,$
+then
+(7) $d_{1}\leq\dim_{H}(K(B))\leq d_{2}.$
+Let us discuss how to find estimates for $d_{1}$ and $d_{2}$. The iterates of the Gauss map are given explicitly by
+$\psi|_{I_{j}}(x)=\dfrac{q^{(j)}_{r_{j}}x-p^{(j)}_{r_{j}}}{p^{(j)}_{r_{j}-1}-q^{(j)}_{r_{j}-1}x}$
+where $\dfrac{p^{(j)}_{k}}{q^{(j)}_{k}}=[0;b^{(j)}_{1},\dots,b^{(j)}_{k}$] and $\beta_{j}=(b^{(j)}_{1},\dots,b^{(j)}_{r_{j}})$.
+Hence
+$(\psi|_{I_{j}})^{\prime}(x)=\frac{(-1)^{r_{j}-1}}{(p^{(j)}_{r_{j}-1}-q^{(j)}_{r_{j}-1}x)^{2}}.$
+Let $x=[c_{0},c_{1},c_{2},\dots]$ and $\frac{p_{n}}{q_{n}}=[c_{0},c_{1},\dots,c_{n}]$. Then
+$\frac{1}{2q_{n}q_{n+1}}<\frac{1}{q_{n}(q_{n}+q_{n+1})}<\left|x-\frac{p_{n}}{q_{n}}\right|<\frac{1}{q_{n}q_{n+1}},$
+and therefore
+$\frac{1}{2q_{n+1}}<|q_{n}x-p_{n}|<\frac{1}{q_{n+1}}.$
+Therefore,
+$(q_{r_{j}}^{(j)})^{2}<|(\psi|_{I_{j}})^{\prime}(x)|=\frac{1}{(p^{(j)}_{r_{j}-1}-q^{(j)}_{r_{j}-1}x)^{2}}<(2q_{r_{j}}^{(j)})^{2}.$
+Thus
+$(q_{r_{j}}^{(j)})^{2}\leq\lambda_{j}=\inf|\psi^{\prime}|_{I_{j}}|\leq\Lambda_{j}=\sup|\psi^{\prime}|_{I_{j}}|\leq(2q_{r_{j}}^{(j)})^{2}.$
+Consider the set $P_{n}(x,m)=\{(a,b)\in Q_{n}(x):x/2^{m/2}<b\leq x/2^{(m-1)/2}\}$.
+Notice that writing $a/b=[0;b_{1},\dots,b_{\ell}]$, each $0;b_{1},\dots,b_{i}]$ belongs to a different $P_{n}(x,m)$ for $1\leq b_{i}\leq\ell$.
+In particular the intervals $I(b_{1},\dots,b_{\ell})$ where $(a,b)\in P_{n}(x,m)$ form a Markov partition of the Gauss-Cantor set $C(n)$.
+By (7) we have that
+$1=\sum_{i=1}^{|P_{n}(x,m)|}\lambda_{j}^{-d_{2}}\leq|P_{n}(x,m)|x^{-2d_{2}}2^{d_{1}m}\leq|P_{n}(x,m)|x^{-2\dim_{H}(C(n))}2^{m},$
+whence
+$|Q_{n}(x)|=\sum_{m\geq 1}|P_{n}(x,m)|\geq x^{2\dim_{H}(C(n))}.$
+Similarly
+$1=\sum_{i=1}^{|P_{n}(x,m)|}\Lambda_{j}^{-d_{1}}\geq|P_{n}(x,m)|x^{-2d_{1}}2^{d_{2}(m+1)}\geq|P_{n}(x,m)|x^{-2\dim_{H}(C(n))}2^{m+1},$
+whence whence
+$|Q_{n}(x)|=\sum_{m\geq 1}|P_{n}(x,m)|\leq 2x^{2\dim_{H}(C(n))}.$
+2.2. The subshift $\Sigma_{3}$
 Given $t>0$, define the subshift
-$\Sigma_{t}=\left\{\underline{a}\in(\mathbb{N}^{*})^{\mathbb{Z}}\ \mid\ \sup_{n\in\mathbb{Z}}\lambda_{n}(\underline{a})\leq t\right\}.$
+$\Sigma_{t}=\left\{\underline{a}\in(\mathbb{N}_{>0})^{\mathbb{Z}}\ \mid\ \sup_{n\in\mathbb{Z}}\lambda_{n}(\underline{a})\leq t\right\}.$
 Note that for $t\in(0,\infty)$, one only needs finitely many symbols because $\Sigma_{t}\subset\{1,\dots,\lfloor t\rfloor\}^{\mathbb{Z}}$.
@@
 However, the set $K_{3}$ is a Cantor set, as we will see.
-The subshift $\Sigma_{3}$ was characterized in [10]: after substituting $22\mapsto a$ and $11\mapsto b$ it coincides with the subshift of all Sturmian bi-infinite words (eventually periodic and aperiodic).
+The subshift $\Sigma_{3}$ was characterized in [14]: after substituting $22\mapsto a$ and $11\mapsto b$ it coincides with the subshift of all balanced bi-infinite words (eventually periodic and aperiodic).
 It is well known that the closure of any aperiodic Sturmian bi-infinite word gives a minimal infinite subshift, so in particular, any orbit in that closure is an accumulation point.
-On the other hand, the eventually periodic Sturmian bi-infinite words can be approximated by the aperiodic (in the shift topology), thus any orbit of $\Sigma_{3}$ is an accumulation point.
+On the other hand, the eventually periodic balanced bi-infinite words can be approximated by the aperiodic (in the shift topology), thus any orbit of $\Sigma_{3}$ is an accumulation point.
 In particular the shift $\Sigma_{3}$ is homeomorphic to a Cantor space and consequently the set $K_{3}$ is a Cantor subset of $\mathbb{R}$.
-We define $\Sigma(t,n)$ to be the set of length-$n$ subwords of sequences in $\Sigma_{t}$. From [4, Theorem 1.1] we have the following:
+The shift $\Sigma_{3}$ can be described via the SS–adic representation of balanced sequences.
+Background material and proofs of the statements below can be found in [15].
+We call the substitutions $U$ and $V$ the inner renormalization operators, defined by
+[…]
+Let $(\alpha,\beta)$ be a pair of finite words over the alphabet $\{a,b\}$. Define the exterior renormalization operators $\overline{U},\overline{V}$ by
+$\overline{U}:(\alpha,\beta)\mapsto(\alpha\beta,\beta),\quad\overline{V}:(\alpha,\beta)\mapsto(\alpha,\alpha\beta).$
+The Cohn’s tree is a binary tree of finite words over $\{a,b\}$ defined as follows: take as root the word $ab$ and given a vertex, its children in the next layer are obtained by applying $U$ and $V$.
+Let $P$ be the set of vertices of Cohn’s tree.
+These finite words together with $\{a,b\}$ are called (lower) Christoffel words.
+It is a fact that a subword is subfactor of a balanced sequence if and only if is subfactor of some Christoffel word if and only if is a subfactor of some Sturmian sequence.
+By [5, Lemma 3.11] we have a control of how large is this Christoffel word.
+Lemma 2.2.
+Let $w$ be a factor of a word in $P$. Then the length of the shortest word in $P$ containing $w$ is strictly smaller than $3|w|$.
+Analogously, one can define a tree of alphabets from word pairs as follows: the root is $(a,b)$ and given a vertex, its children in the next layer are obtained by applying $\overline{U}$ and $\overline{V}$.
+The set of vertices of this tree $\overline{P}$ are also called Christoffel pairs.
+The Cohn’s tree and the tree of alphabets are related through the concatenation operation: given an alphabet $(\alpha,\beta)\in\overline{P}$ we have that $\alpha\beta\in P$ and conversely, given any Christoffel word $w\in P$, there is a unique factorization $w=\alpha\beta$ where $(\alpha,\beta)\in\overline{P}$.
+Moreover, given any $(\alpha,\beta)\in\overline{P}$, there are some substitutions $R_{1},\dots,R_{n}$ such that $\alpha=(R_{1}\dots R_{n})(a)$ and $\beta=(R_{1}\dots R_{n})(b)$.
+Another useful property of lower Christoffel words $P$ is that they are completely determined by counting their letters.
+More precisely, for $w\in P$, the Frobenius coordinates $(|w|_{a}+|w|_{b},|w|_{b})$ uniquely determine $w$, where $|w|_{c}$ denotes the number of occurrences of the letter $c\in\{a,b\}$.
+Furthermore, every word in $P$ is primitive, that is, no $w\in P$ is a proper power of a shorter word.
+Consequently, all cyclic permutations of a given $w$ are distinct.
+In addition, cyclic permutations of two distinct words $w,w^{\prime}\in P$ cannot coincide, since their Frobenius coordinates are different.
+Continued fractions are closely related to the Christoffel pairs.
+Indeed, for each continued fraction $0;a_{1},\dots,a_{n}]$ we can associate a unique Christoffel pair as follows.
+Define $(\alpha_{0},\beta_{0})=(a,b)$, $(\alpha_{1},\beta_{1})=\overline{V}^{a_{1}-1}(\alpha_{0},\beta_{0})$, $(\alpha_{2},\beta_{2})=\overline{U}^{a_{2}}(\alpha_{1},\beta_{1})$, $(\alpha_{3},\beta_{3})=\overline{V}^{a_{3}}(\alpha_{2},\beta_{2})$ and so on.
+In particular we will have that $q_{0}=|\alpha_{0}|/2$, $q_{1}=|\beta_{1}|/2$, $q_{2}=|\alpha_{2}|/2$, $q_{3}=|\beta_{3}|/3$ and so on, where $q_{1},\dots,q_{n}$ are the denominators of the convergents of $0;a_{1},\dots,a_{n}]$.
+This process can be imagined as the mediant process used to build the convergents of the continued fraction $0;a_{1},\dots,a_{n}]$.
+$\alpha$$\beta$$\alpha\beta$$\alpha\beta^{2}$$\alpha\beta^{3}$$\alpha\beta^{4}\dots$
+$\alpha$$\beta$$\alpha\beta^{3}\dots$$\alpha\beta^{3}((\alpha\beta^{3})^{2}\beta)^{4}$$(\alpha\beta^{3})^{2}\beta$
+Figure 1. Example of how constructing alphabets is similar to build the convergents of a number $[0;1,3,2,4]$.
+Recall that given fractions $0<\frac{p}{q}<\frac{r}{s}$ the mediant is
+$\frac{p}{q}\oplus\frac{r}{s}=\frac{p+r}{q+s}.$
+Given $n\geq 1$, the Farey sequence $\mathcal{F}_{n}$ of order $n$ is the set of reduced coprime fractions $0\leq\frac{p}{q}\leq 1$ with $1\leq q\leq n$ increasingly ordered.
+Two fractions $\frac{p}{q}<\frac{r}{s}$ are Farey neighbors in some $\mathcal{F}_{n}$ if and only if psqr=1ps-qr=1 and $\max\{q,s\}\leq n<q+s$.
+The mediant $\frac{p+r}{q+s}$ will appear between $\frac{p}{q}$ and $\frac{r}{s}$ precisely at term $\mathcal{F}_{q+s}$.
+The rule used to construct the tree of alphabets is related to the mediant rule used in the construction of the Farey sequence.
+We can build a bijection $\Theta$ from the Christoffel words $P\cup\{a,b\}$ to the positive fractions $\mathbb{Q}\cap[0,1]$ as follows.
+Given $w\in P\cup\{a,b\}$, we define $\Theta(w)=\frac{|w|_{b}}{|w|_{a,b}}$ where $|w|_{a,b}$ is the length of $w$ over the alphabet $\{a,b\}$ and $|w|_{b}$ is the number of letters $b$ in $w$.
+The following fact is consequence of [17, Lemma 3.1].
+Lemma 2.3.
+For each Farey neighbors $\frac{p}{q}<\frac{r}{s}$, there is a unique pair of words $(\alpha,\beta)\in\overline{P}$ such that $\Theta(\alpha)=p/q$ and $\Theta(\beta)=r/s$.
+We define $\Sigma(t,n)$ to be the set of length-$n$ subwords of sequences in $\Sigma_{t}$. From [5, Theorem 1.1] we have the following:
 Theorem 2.4.
@@
 Remark 2.6.
-In fact, it is possible to prove that $|\Sigma(3,n)|\leq 9n^{3}$ for all $n\geq 1$, see [4, Corollary 3.13].
+In fact, it is possible to prove that $|\Sigma(3,n)|\leq 9n^{3}$ for all $n\geq 1$, see [5, Corollary 3.13].
 Lemma 2.7.
@@
 In view of this, we introduce the following lemma concerning the fractal geometry of Cantor sets.
-The next lemma is a slightly modified version of [5, Lemma 1].
+The next lemma is a slightly modified version of [8, Lemma 1].
 The proof is similar, and we include it here for completeness.
@@
 In particular, this shows that $k^{-1}(3)$ can be covered by the countable union
-$k^{-1}(3)\subset\bigcup_{n_{0}\geq 2}\bigcup_{(a_{1},\dots,a_{n_{0}-1})\in(\mathbb{N}^{*})^{n_{0}-1}}\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}.$
+$k^{-1}(3)\subset\bigcup_{n_{0}\geq 2}\bigcup_{(a_{1},\dots,a_{n_{0}-1})\in(\mathbb{N}_{>0})^{n_{0}-1}}\{[0;a_{1},\dots,a_{n_{0}-1},\gamma]:\gamma\in K_{3+6^{-m}}\}.$
 where $K_{t}$ was defined on (1).
@@
 Here, $\Gamma_{h}(r)=r\cdot\inf_{0<s\leq r}\frac{h(s)}{s}$ and $\mathbb{L}$ denotes the set of Liouville numbers.
-It is proved in [9, Lemma 2.2] that for all dimension functions $h$ and all sets $E\subset\mathbb{R}$, we have:
+It is proved in [12, Lemma 2.2] that for all dimension functions $h$ and all sets $E\subset\mathbb{R}$, we have:
 $\mathcal{H}^{\Gamma_{h}}(E)\leq\mathcal{H}^{h}(E)\leq 2\mathcal{H}^{\Gamma_{h}}(E).$
@@
 has infinitely many solutions $p/q\in\mathbb{Q}$. By the definition of the function $k$, we have that $k(x+1/2)=k(y)\geq 4$ (in fact $k(y)\geq 12$).
-4. The set $K_{3}$
+4. Dimensionality of $K_{3}$
+For the Cantor set
+$K_{3}=\{[0;a_{1},\dots,a_{n},\dots]\ \mid\ \text{there exists $(a_{-n})_{n\geq 0}\in(\mathbb{N}_{>0})^{\mathbb{N}}$ such that}\\ [a_{k};a_{k+1},\dots,]+[0;a_{k-1},a_{k-2},\dots]\leq 3,\forall k\in\mathbb{Z}\},$
+we obtain a reasonably explicit description of its generalized Hausdorff measures. However, due to additional combinatorial complications, the result is not as precise as in the case of $k^{-1}(3)$.
+Recall that for any fixed $n\geq 1$, the set $\{I(w):w=(a_{1},\dots,a_{n})\in\Sigma(3,n)\}$ is a covering of $K_{3}$ by closed intervals.
 Lemma 4.1.
@@
 (16) $h(\varepsilon_{k})\leq\frac{1}{|\log\varepsilon_{k}|^{3+d}},\quad\text{for all $k$}.$
-Given $k$, let $n_{k}=\lceil\log_{2}\varepsilon_{k}\rfloor$.
-Note that for any $n_{k}\geq 1$, the set $\{I(w):w=(a_{1},\dots,a_{n_{k}})\in\Sigma(3,n_{k})\}$ is a covering of $K_{3}$ by closed intervals.
-Since $\operatorname{diam}I(w)<2^{-n_{k}}$, we have that
-$\sum_{w\in\Sigma(3,n_{k})}h(\operatorname{diam}I(w))$ $\leq|\Sigma(3,n_{k})|h(\varepsilon_{k})$
+For a given $k$, define $n_{k}=\lceil\log_{2}\varepsilon_{k}\rfloor$. Taking $w=(a_{1},\dots,a_{n_{k}})\in\Sigma(3,n_{k})$, since $\operatorname{diam}I(w)<2^{-n_{k}}$, we have that
+$\sum_{w\in\Sigma(3,n_{k})}h(\operatorname{diam}I(w))$ $\leq|\Sigma(3,n_{k})|\cdot h(\varepsilon_{k})$
 $\leq\frac{9n_{k}^{3}}{|\log\varepsilon_{k}|^{3+d}}\leq\frac{9n_{k}^{3}}{(\log(2)-1)^{3+d}n_{k}^{3+d}}=\frac{9}{(\log(2)-1)^{3+d}n_{k}^{d}}.$
 Since $n_{k}\to\infty$ as $k\to\infty$, this proves the lemma. ∎
+We can actually get a more precise estimation of the gauge functions that give zero measure.
+For example we can prove that the function $h(\varepsilon)=|\log\varepsilon|^{-3}$ gives $\mathcal{H}^{h}(K_{3})=0$.
+[…]
+Lemma 4.2.
+For the dimension function
+$g(\varepsilon)=\frac{\log|\log\varepsilon|}{|\log\varepsilon|^{3}(\log\log|\log\varepsilon|)},$
+we have that $\mathcal{H}^{g}(K_{3})<\infty$.
+Proof.
+We will consider initially the covering of $K_{3}$ given by the intervals $I(a_{1},\dots,a_{N})$ where $(a_{1},\dots,a_{N})\in\Sigma(3,N)$.
+Since some of these intervals are very small, we will show that they can be covered more efficiently by bigger intervals $I(a_{1},\dots,a_{\ell})$ associated with shorter prefixes.
+Given any word $(a_{1},\dots,a_{N})\in\Sigma(3,N)$, we know by Lemma 2.2 that there is a word $w\in P$ with $N\leq|w|\leq 3N$ such that $(a_{1},\dots,a_{N})$ is a subword of $w$.
+In particular, if $w\neq b$, we can write $w=\alpha$ for some $(\alpha,\beta)\in\overline{P}$ with $|\alpha|>|\beta|$.
+Writing $w=w_{1}\dots w_{|w|}\in\{1,2\}^{|w|}$, since $\alpha\alpha$ is a subfactor of $\alpha^{\infty}\in\Sigma_{3}$, all cyclic permutations $w_{i}\dots w_{|w|}w_{1}\dots w_{i-1}$, $1\leq i\leq|w|$ appear in $\Sigma(3,|w|)$ and thus appear as prefixes of continued fractions in $K_{3}$.
+In particular we have that $a_{1}\dots a_{N}$ is a prefix of $w_{i}\dots w_{|w|}w_{1}\dots w_{i-1}$ for some $1\leq i\leq|w|$.
+We will extend each $(a_{1},\dots,a_{N})$ so that it coincides with that permutation $(a_{1},\dots,a_{N^{\prime}})=(w_{i},\dots,w_{|w|},w_{1},\dots,w_{i-1})$ for some $N^{\prime}\leq 3N$.
+We will refine further this covering by considering all $\alpha\in P$ with $N\leq N^{\prime}=|\alpha|\leq 6N$.
+On the other hand, as explained in Section 2.2, there is a unique continued fraction $0;b_{1},\dots,b_{m}]$ that codifies the Christoffel pair $(\alpha,\beta)\in\overline{P}$ and satisfies $q_{m}=|\alpha|/2$.
+Notice that the denominator of such continued fraction is between $N/2$ and $3N$.
+Let $M$ be a large positive value that depends on $N$ that we will chose later.
+We have two different situations, either there is a coefficient $b_{j+1}>M$ with $j$ minimal or $b_{j}\leq M$ for all $1\leq j\leq m$.
+Now we will consider the first situation.
+Let us count how many such continued fractions with denominator between $N/2$ and $3N$ and such that $M<b_{j+1}\leq 2M$ is the first coefficient larger than $M$ exist.
+Writing $0;b_{1},\dots,b_{j},M]=\frac{p_{r}}{q_{r}}$, any such fraction is of the form $0;b_{1},\dots,b_{j},M+s,\frac{p}{q}]$ where $0<s\leq M$ and $0\leq p/q\leq 1$.
+Writing $0;b_{1},\dots,b_{j},M+s]=\frac{p_{r}+sp_{r-1}}{q_{r}+sq_{r-1}}$, we have that
+$0;b_{1},\dots,b_{j},M+s,q/p]=\frac{q(p_{r}+sp_{r-1})+pp_{r-1}}{q(q_{r}+sq_{r-1})+pq_{r-1}}.$
+Since $q(q_{r}+sq_{r-1})+pq_{r-1}\leq q(q_{r}+Mq_{r-1})+qq_{r-1}<3qq_{r}$, we see that any choice of $\frac{N}{2q_{r}}<q<\frac{N}{q_{r}}$, $0\leq p\leq q$ and $0<s\leq M$, gives $q(q_{r}+sq_{r-1})+pq_{r-1}\in[N/2,3N]$.
+Similarly we have $\frac{N}{6q_{r}}<q<\frac{3N}{q_{r}}$.
+Therefore we have at least $(\frac{N}{2q_{r}})^{2}$ and at most $\left(\frac{3N}{q_{r}}\right)^{2}$ choices for $0\leq p\leq q$, and we have $M$ choices for $s$.
+In resume, we have between $\frac{N^{2}M}{4q_{r}^{2}}$ and $\frac{9N^{2}M}{q_{r}^{2}}$ such continued fractions $0;b_{1},\dots,b_{m}]=[0;b_{1},\dots,b_{j},M+s,\frac{p}{q}]$.
+Now, since permutations of $\alpha$ also give different possibilities for $(a_{1},\dots,a_{N^{\prime}})$, we have between $\frac{N^{3}M}{4q_{r}^{2}}$ and $\frac{54N^{3}M}{q_{r}^{2}}$ intervals $I(a_{1},\dots,a_{N^{\prime}})$ with $(a_{1},\dots,a_{N^{\prime}})$ a cyclic permutation of $\alpha$ associated to a unique $0;b_{1},\dots,b_{j},M+s,q/p]$.
+All of them have diameter at most $\left(\frac{3+\sqrt{5}}{2}\right)^{-6N+1}$.
+On the other hand we know that $|\Sigma(3,N^{\prime})|\leq 9(N^{\prime})^{3}\leq 1944N^{3}$, so if we denote
+$\mathcal{C}=\left\{(p_{r},q_{r}):\begin{subarray}{c}\displaystyle 1\leq p_{r}\leq q_{r}\leq 3N,\gcd(p_{r},q_{r})=1,\\ \displaystyle p_{r}/q_{r}=[0;b_{1},\dots,b_{j},M],1\leq b_{i}\leq M,1\leq i\leq j\end{subarray}\right\},$
+we have summing over all such intervals
+(17) $\sum_{(p_{r},q_{r})\in\mathcal{C}}\frac{1}{q_{r}^{2}}\leq\frac{|\Sigma(3,6N)|}{54N^{3}M}\leq\frac{36}{M}.$
+[…]
+(18) $\sum_{(p_{r},q_{r})\in\mathcal{C}}2q_{r}g\left(\left(\frac{3+\sqrt{5}}{2}\right)^{-2q_{r}+1}\right)\leq C^{\prime}\sum_{(p_{r},q_{r})\in\mathcal{F}}\frac{1}{q_{r}^{2}}\frac{\log q_{r}}{\log\log q_{r}},$
+for some constant $C^{\prime}>0$. Now we will chose $M$ to be equal to $M=C(\log N)/(\log\log N)$ for some absolute constant $C>0$ we will chose later. In particular, using (17) and (18)
+$\sum_{(p_{r},q_{r})\in\mathcal{C}}2q_{r}g\left(\left(\frac{3+\sqrt{5}}{2}\right)^{-q_{r}+1}\right)\leq\frac{C^{\prime}}{C}\sum_{(p_{r},q_{r})\in\mathcal{F}}\frac{M}{q_{r}^{2}}=\frac{36C^{\prime}}{C}<\infty.$
+If $\min\{|\hat{\alpha}|,|\hat{\beta}|\}=2$, then $j+1=1$ and $(\hat{\alpha},\hat{\beta})=(a,b)$ and $(a_{1},\dots,a_{N^{\prime}})$ has no prefix that is permutation of $\tilde{\theta}\tilde{\theta}$, it is because either $2b^{M+1}$, $b^{M+1}$ or $a^{M+1}$ is a prefix of $(a_{1},\dots,a_{N^{\prime}})$.
+Since the diameter of the intervals $I(2b^{M+1})$, $I(b^{M+1})$, $I(a^{M+1})$ goes to zero provided $M$ goes to infinity as $N$ goes to infinity, the contribution of these intervals can be ignored.
+In conclusion the contribution of all the intervals in the first situation is bounded by some absolute constant.
+On the other hand, in the second situation, if all coefficients are $b_{j}\leq M$ for $1\leq j\leq m$, then we will not change $I(a_{1},\dots,a_{N^{\prime}})$.
+By (6) we have at most $2(3N)^{2\dim_{H}(C(M))}$ continued fractions $0;b_{1},\dots,b_{m}]$ with denominator at most $3N$ and $1\leq b_{i}\leq M$ for $1\leq i\leq m$.
+For each such continued fraction, it corresponds a unique $\alpha\in P$ and at most $|\alpha|\leq 6N$ permutations.
+Using Jarnik’s estimate there is an absolute constant $C_{1}>0$ such that $\dim_{H}(C(M))<1-C_{1}/M$, therefore we have at most $4(3N)^{2\dim_{H}(C(M))+1}\leq 108N^{3}N^{-2C_{1}/M}$ intervals.
+Now we will fix $C=2C_{1}$, so we have at most $108N^{3}/\log N$ such intervals with diameter at most $\left(\frac{3+\sqrt{5}}{2}\right)^{-6N+1}$.
+Therefore, they have gg–measure at most
+$\frac{108N^{3}}{\log N}g\left(\left(\frac{3+\sqrt{5}}{2}\right)^{-6N+1}\right)\leq\hat{C}\frac{1}{\log\log N},$
+for some absolute constant $\hat{C}>0$. This shows that the contribution of these intervals goes to zero.
+Lemma 4.3.
+For a large enough constant $\hat{C}>0$, the dimension function
+$g(\varepsilon)=\frac{\exp(\hat{C}\sqrt{\log|\log\varepsilon|})}{|\log\varepsilon|^{3}}$
+satisfies $\mathcal{H}^{g}(K_{3})=\infty$. In particular for any dimension function $h$ such that
+$\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}<3,$
+we have $\mathcal{H}^{h}(K_{3})=\infty$.
+Proof.
+We will use Lemma 3.3.
+We chose $C_{n}=\bigcup\{I(w):w=(a_{1},\dots,a_{n})\in\Sigma(3,2^{n^{2}})\}$ to be our coverings so in particular $K_{3}=\bigcap_{n\geq 1}C_{n}$.
+We claim that there is a constant $C>0$ such that any interval $I(w)$ with $w\in\Sigma(3,2^{n^{2}})$ is subdivided in at least $C\cdot 2^{3\cdot(2n+1)}$ intervals I(w~)I($\tilde${w}) with $\tilde{w}\in\Sigma(3,2^{(n+1)^{2}})$, or equivalently, each $w\in\Sigma(3,2^{n^{2}})$ has at least $C\cdot 2^{3(2n+1)}$ different such continuations $\tilde{w}\in\Sigma(3,2^{(n+1)^{2}})$.
+Now we will prove the claim.
+Given any word $w\in\Sigma(3,n^{\prime})$, by Lemma 2.2 there is a Christoffel word $\alpha\in P\cup\{a,b\}$ with $2m=|\alpha|<3n^{\prime}$ such that $w$ is subfactor of $\alpha$.
+By Lemma 2.3, there is another Christoffel word $\beta\in P\cup\{a,b\}$ such that $(\alpha,\beta)\in\overline{P}$ is an alphabet and such that $\Theta(\alpha)<\Theta(\beta)$ are Farey neighbors in $\mathcal{F}_{m}$.
+In particular, $\max\{|\alpha|_{a,b},|\beta|_{a,b}\}\leq m<|\alpha|_{a,b}+|\beta|_{a,b}$.
+Denote by $L_{k}$ the set of all words $l_{1}\dots l_{k}\in\{a,b\}^{k}$ that appear as subfactors of some word in $\Sigma_{3}$.
+It is known [14] that $L_{k}$ is the same as the Sturmian factors of length $k$ and it is known [9] that we have $|L_{k}|\sim k^{3}/\pi^{2}$.
+For our purposes, it suffices to prove that there is a constant $\tilde{C}>0$ such that $|L_{k}|\geq 2\tilde{C}k^{3}$ for all $k$.
+Let us give an alternative proof of this fact that it is useful for the next section.
+Recall that given a Christoffel word $w\in P\cup\{a,b\}$ we can associate a unique Farey fraction $\Theta(w)=\frac{|w|_{b}}{|w|_{a,b}}$ where $|w|_{b}$ denotes the number of letters $b$ in $w$ and $|w|_{a,b}$ is the length of $w$ over the alphabet $\{a,b\}$.
+In particular if $\mathcal{F}_{k}$ denotes the kk–term of the Farey sequence, we can find at least $|\mathcal{F}_{k}|\sim\frac{3k^{2}}{\pi}$ elements in $L_{k}$.
+Since cyclic permutations also appear as subfactors of longer Christoffel words and since $|\mathcal{F}_{k}|=|\mathcal{F}_{k-1}|+\varphi(k)$ where $\varphi$ denotes the Euler function, we have the known formula $|L_{k}|=1+\sum_{i=1}^{k}(k-i+1)\varphi(k)\sim\frac{k^{3}}{\pi^{2}}$.
+On the other hand, since the operation of exchanging $a$ with $b$ preserves factors in $L_{k}$, precisely half of the words in $L_{k}$ begin with $a$, say $\{\eta_{1},\dots,\eta_{|L_{k}|/2}\}$.
+Since $(\alpha,\beta)\in\overline{P}$, there are some inner renormalization operators $R_{1},\dots,R_{\ell}\in\{U,V\}$ such that if $\alpha=R(a)$ and $\beta=R(b)$ where $R=R_{1}\dotsb R_{\ell}$.
+In particular, each $R(\eta_{j})$, $1\leq j\leq|L_{k}|/2$ is also a Sturmian factor of length over the alphabet $\{1,2\}$ of at most $mk<\frac{3}{2}n^{\prime}k$.
+Each such factor $R(\eta_{j})$ begins with $\alpha$ and since $w$ is a subfactor of $\alpha$, each $R(\eta_{j})$ provides a continuation of $w$ of length over the alphabet $\{1,2\}$ at most $\frac{3}{2}n^{\prime}k$.
+Since any Sturmian factor can be extended arbitrarily, we conclude that any $w\in\Sigma(3,n^{\prime})$ has at least $|L_{k}|/2\geq\tilde{C}k^{3}$ continuations $\tilde{w}\in\Sigma(3,3n^{\prime}k)$.
+Consequently, to finish the proof of the claim, we chose $k=\lfloor 2^{2n+2}/3\rfloor$ and $n^{\prime}=2^{n^{2}}$.
+Notice that by (4) the size of each interval of $C_{n}$ is at least $\varepsilon_{n}=(3+2\sqrt{2})^{-2^{n^{2}}-1}$.
+According to the notation of Lemma 3.3, we will have $N_{n}=\prod_{j<n}(C\cdot 2^{3(2j+1)})=C^{n-1}2^{3n^{2}-3}$.
+In particular, for a large enough constant $\hat{C}>0$, the dimension function
+$g(\varepsilon)=\frac{\exp(\hat{C}\sqrt{\log|\log\varepsilon|})}{|\log\varepsilon|^{3}}$
+satisfies
+$\lim_{n\to\infty}N_{n}\cdot g(\varepsilon_{n+1})\geq\lim_{n\to\infty}\frac{\exp(n\log(C/64)+\hat{C}(n+1)\sqrt{\log 2})}{64C\log(3+2\sqrt{2})^{3}}=\infty,$
+so by Lemma 3.3 we obtain $\mathcal{H}^{g}(K_{3})=\infty$. Moreover, the function $g$ clearly satisfies
+$\lim_{\varepsilon\to 0}\frac{|\log g(\varepsilon)|}{\log|\log\varepsilon|}=\lim_{\varepsilon\to 0}\left|\frac{\hat{C}}{\sqrt{\log|\log\varepsilon|}}-3\right|=3.$
+Finally, by the hypothesis on the dimension function $h$, for all sufficiently small $\varepsilon>0$ we have $h(\varepsilon)>g(\varepsilon)$.
+Consequently $\mathcal{H}^{h}(K_{3})\geq\mathcal{H}^{g}(K_{3})=\infty$.
+∎
+5. Dimensionality of $B_{3}$
+The following is a consequence of the main theorem of [3]. Recall that given a word $w=w_{1}\dots w_{n}\in\{a,b\}^{n}$, we denote $w^{+}=w_{2}\dots w_{n}$ and $w^{T}=w_{n}\dots w_{1}$.
+Theorem 5.1.
+Let $x=[x_{0};x_{1},x_{2},\dots]$ be such that $\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}$ has only finitely many solutions and let $N\in\mathbb{N}=\{0,1,\dots\}$ minimal such that $\lambda_{N}(x)\leq 3$ for all $n\geq N+1$ and $k(x)=3$.
+Then there is a sequence of alphabets $(\alpha_{n+1},\beta_{n+1})\in\{\overline{U}(\alpha_{n},\beta_{n}),\overline{V}(\alpha_{n},\beta_{n})\}$ with both renormalization operators $\overline{U}$ and $\overline{V}$ appearing infinitely many times and such that:
+  • if $x_{N+1}=1$ then $x_{N+1}x_{N+2}x_{N+3}\ldots=\lim_{n\to\infty}\beta_{n}^{T}$ and if $N\geq 1$ then $0;x_{N},\dots,x_{1}]<[0;\lim_{n\to\infty}\alpha_{n}]$; or
+  • if $x_{N+1}=2$ then $x_{N+1}x_{N+2}x_{N+3}\ldots=\lim_{n\to\infty}2\alpha_{n}^{+}$ and if $N\geq 1$ then $0;x_{N},\dots,x_{1}]<[0;2,\lim_{n\to\infty}\beta_{n}^{T}]$.
+Conversely, for any such continued fraction $x=[x_{0};x_{1},x_{2},\dots]$ the inequality $\left|x-\frac{p}{q}\right|<\frac{1}{3q^{2}}$ has at most $N$ solutions.
+Let us denote by $B_{3}(N)$ the set of $x\in B_{3}\cap(0,1)$ such that $N\in\mathbb{N}$ is minimal with $\lambda_{n}(x)\leq 3$ for all $n\geq N+1$. Since
+$B_{3}(N)\subset\bigcup_{(x_{1},\dots,x_{N})\in(\mathbb{N}_{>0})^{N}}\{[0;x_{1},\dots,x_{N},1/\gamma]:\gamma\in B_{3}(0)\cap(0,1)\},$
+we have that $\mathcal{H}^{h}(B_{3}(N))\leq\mathcal{H}^{h}(B_{3}(0))$.
+On the other hand, we clearly have $\mathcal{H}^{h}(B_{3})\geq\mathcal{H}^{h}(B_{3}(0))$.
+Therefore, to estimate the generalized Hausdorff dimension of $B_{3}$, we will consider instead $B_{3}(0)$, that is, the set
+$B_{3}(0)=\left\{x\in\mathbb{R}\setminus\mathbb{Q}:\left|x-\frac{p}{q}\right|\geq\frac{1}{3q^{2}}\text{ for all }\frac{p}{q}\in\mathbb{Q}\right\}.$
+Now another reduction can be made to the study of the dimensionality of the set $B_{3}(0)$.
+Indeed, given a sequence $d_{1},d_{2},\dots$ of positive integers, denote $i_{n}=d_{n}$ and $\overline{R}_{n}=\overline{U}$ if $n$ is even and $i_{n}=d_{n}-1$ and $\overline{R}_{n}=V$ is $n$ is odd, then we have that $(x_{0},x_{1},x_{2},\dots)=\lim_{n\to\infty}\alpha_{n}$ where $(\alpha_{n},\beta_{n})=\left(\overline{R_{n}}^{i_{n}}\dotsb\overline{R_{1}}^{i_{1}}\right)(a,b)$ defines uniquely the number $0;x_{1},x_{2},\dots]$ of $B_{3}(0)$.
+It turns out that if we change $(x_{0},x_{1})=(2,2)$ by $(1,1)$, then $(1,1,x_{2},x_{3},\dots)=\lim_{n\to\infty}\beta_{n}^{T}$, so $0;1,1,x_{2},x_{3},\dots]$ also defines the corresponding number of $B_{3}(0)$.
+To see this, denote $\alpha^{b}=b\alpha^{+}$, so we have that $(1,1,x_{2},\dots)$ begins with $\alpha_{2n}^{b}=(\alpha_{2n-1}^{d_{2n-1}}\beta_{2n-1})^{b}(\alpha_{2n-1}^{d_{2n-1}-1}\beta_{2n-1})^{d_{2n}-1}=(\beta^{T}(\alpha_{2n-1}^{T})^{d_{2n-1}-1}\alpha_{2n-1}^{b})(\alpha_{2n-1}^{d_{2n-1}-1}\beta_{2n-1})^{d_{2n}-1}$, begins with $\beta_{2n-1}^{T}$.
+Since this holds for all $n$, the sequence $(1,1,x_{2},\dots)$ coincides with the limit $\lim_{n\to\infty}\beta_{n}^{T}$.
+Therefore, it suffices to consider numbers of the form $0;2,\lim_{n\to\infty}\alpha_{n}^{+}]$ in $B_{3}(0)$.
+Moreover, since $\frac{1}{2+B_{3}(0)}$ is a Bi-Lipschitz copy of $B_{3}(0)$, it suffices to consider the numbers of the form $0;\lim_{n\to\infty}\alpha_{n}]$ where $\alpha_{n}$ is obtained through a sequence of renormalization operators where both $\overline{U}$ and $\overline{V}$ appear infinitely many times.
+The same proof of Lemma 4.2 but ignoring the permutation of factors give the following.
+Lemma 5.2.
+For the dimension function
+$g(\varepsilon)=\frac{\log|\log\varepsilon|}{|\log\varepsilon|^{2}(\log\log|\log\varepsilon|)},$
+we have that $\mathcal{H}^{g}(B_{3}(0))<\infty$. In particular for any dimension function $h$ such that
+$\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}>2,$
+we have $\mathcal{H}^{h}(B_{3})=0$.
+Similarly, the same proof of Lemma 4.3 but ignoring the permutation of factors give the following.
+Lemma 5.3.
+For a large enough constant $\tilde{C}>0$, the dimension function
+$g(\varepsilon)=\frac{\exp(\tilde{C}\sqrt{\log|\log\varepsilon|})}{|\log\varepsilon|^{2}}$
+satisfies $\mathcal{H}^{g}(B_{3}(0))=\infty$. In particular for any dimension function $h$ such that
+$\limsup_{\varepsilon\to 0}\frac{|\log h(\varepsilon)|}{\log|\log\varepsilon|}<2,$
+we have $\mathcal{H}^{h}(B_{3})=\infty$.
 Acknowledgements